### v1

# Radar-Guided Polynomial Fitting for Metric Depth Estimation

Abstract
We propose PolyRad, a novel radar-guided depth estimation method that introduces polynomial fitting to transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps. Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a simple yet fundamental insight: using polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges. Although MDE models often infer reasonably accurate local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale-and-shift transformation insufficient given three or more of these regions. In contrast, PolyRad generalizes beyond linear transformations and is able to correct such misalignments by introducing inflection points. Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces monotonicity via first-derivative regularization. PolyRad achieves state-of-the-art performance on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets, outperforming existing methods by 30.3% in MAE and 37.2% in RMSE.
1 Introduction
Metric 3D reconstruction is critical for spatial tasks such as self-driving [34, 14], where it is necessary for one to perceive the structure of the 3D environment in order to navigate. In many such systems, multiple sensors—including cameras, lidar, and radar provide complementary information about the 3D scene. While lidar sensors offer dense, accurate point clouds, they are expensive and not widely available [39]. In contrast, radar sensors, particularly millimeter wave (mmWave) radars [20], return only about a hundred points per frame and are noisier [15]. Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles [10].
Figure 1: Once an MDE model predicts incorrect relative depths between three or more objects, a linear scale-and-shift (dashed) cannot resolve this misalignment. Instead, we learn a polynomial fit with inflection points, enabling flexible non-uniform corrections.
Unless camera baselines are known, images offer scaleless reconstruction. However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint. Training robust models that can generalize across these factors often demands large-scale datasets, which are costly and time-intensive to collect. As a result, leveraging pretrained monocular depth estimation (MDE) models [41, 56, 4, 37] has emerged as a practical alternative. However, as these models only take in a single image, where reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or metric depth that lacks the fidelity needed (see Tab. 2) for applications requiring accurate metric-scale reconstruction, e.g., navigation.
Existing approaches [60, 51, 58, 18, 8, 59] attempt to transform these MDE predictions into accurate metric depth maps using a global linear (scale-and-shift) transformation. This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene. While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other. Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (see Fig. 1). Empirically, a learned linear fit on a challenging dataset such as nuScenes performs worse than even a deterministic estimator like median scaling (see Tab. 2), indicating susceptibility to outliers.
In this work, we challenge the common misconception that scale ambiguity in MDE is only up to an unknown global scale and shift. Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, thereby enabling more flexible non-uniform adjustments. Our approach, PolyRad, transforms the predictions of pretrained MDE models by: learning prototypical patterns in the configuration of radar points, establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and predicting polynomial coefficients that adaptively fit scaleless depth into metric depth. This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and inter-region misalignments. While one may adjust every prediction (pixel) of an MDE model [25], the number of degrees of freedom is the total number of pixels. Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem. On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular. To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models, making our approach a novel solution not only for radar-guided depth estimation but also for any task that leverages foundation models and refines their predictions using additional inputs as guidance.
However, learning such a flexible transformation is challenging because higher-order polynomials introduce a vast function space with many free parameters. Unlike a simple linear fit that relies on only scale and bias terms, polynomial fitting allows for numerous inflection points and nonlinearities. This expressiveness can inadvertently lead to harmful non-monotonic transformations—where the relative depth orderings of points are incorrectly reversed—if not properly constrained. We address this by introducing a regularization term that enforces the first derivative of the predicted metric depth with respect to the input MDE depth to generally remain close to 1, ensuring that incremental changes in the input MDE depth result in approximately equal changes in the predicted depth within local regions where MDE reconstructs structure with high fidelity. In essence, this regularizer encourages a smooth, monotonic transformation, mitigating unstable oscillations that can arise from overfitting high-degree polynomials while still enabling necessary corrections for cross-region misalignments.
Our contributions: We propose (1) a novel radar-guided depth estimation method that leverages polynomial fitting to transform scaleless MDE predictions into accurate metric depth maps. Our approach (2) introduces a fundamental insight: using polynomial coefficients predicted from radar points to enable non-uniform, depth-dependent corrections. We present (3) a principled geometric approach, where polynomial transformations naturally introduce inflection points that can correct misalignments between local areas that a global scale-and-shift transformation cannot. We design (4) a novel training objective that encourages monotonicity through first-derivative as regularization, ensuring structural consistency while allowing inter-region depth adjustments. Finally, (5) extensive experiments on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets demonstrate that PolyRad achieves the state of the art, outperforming existing methods by 11.3%, 54.6%, and 35.2%, respectively.
2 Related Work
2.1 Monocular Depth Estimation
With recent advances in the scalability of neural networks [9, 5], monocular depth estimation (MDE) models can infer relative depth from a single image across diverse, unseen domains [22, 40, 55, 56, 37, 4], benefiting from large-scale datasets. Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation [16, 2], uncertainty estimation [38], optical flow [62], and visual odometry [46], to improve relative depth estimation. However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem. Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity [51]. One way to address this limitation is to incorporate range-sensing modalities such as lidar [21, 11], radar [45], or visual-inertial odometry [52].
2.2 Range-Camera Depth Estimation
Metric-scale depth estimation often requires recovering a dense depth map from sparse range measurements, guided by complementary camera inputs. Image-guided depth completion, such as lidar-camera depth estimation, leverages the strengths of each modality: images offer dense visual context and structural priors, while sparse depth provides absolute metric scale to resolve depth ambiguity. Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers [32, 33], late fusion, which maintains independent branches for different sensors before merging outputs [57, 54, 50, 12, 43], and multi-scale fusion, which captures both local details and global scene structure [23]. U-Net-like architectures with CSPN-based refinement [6] have been widely used for coarse-to-fine depth completion [17, 27, 36], with further improvements from adaptive convolutions [7], deformable convolutions [36, 53], and transformer-based attention mechanisms [42, 61].
2.3 Radar-Camera Depth Estimation
While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions [39]. In contrast, radar provides a cost-effective alternative [19], offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle [35, 47, 44]. The ubiquity of radar sensors [10] in existing automotive and robotic platforms further supports their integration into depth estimation pipelines. Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity [45] of radar point clouds. [26] uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map. [29] leverages Doppler velocity and optical flow to associate radar points with image pixels. Other methods refine radar depth using neighborhood pixel associations [30], height-extension of radar points to address elevation ambiguity [28], or by extracting self-supervisory signals from the radar point cloud [13]. [45] introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion. [24] mitigates distribution artifacts using sparse supervision, while [48] employs a two-stage confidence-driven approach. [49] uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features. Like our method, [25] starts with an initial MDE prediction but refines it through quasi-dense scale estimation, followed by local scale adjustment, rather than through a learned scene-fitting approach.
While these existing methods employ complex architectures to refine radar-derived spatial information, they all operate within the paradigm of completion or confidence-driven fusion. They often rely on multi-stage training, explicit radar-pixel association learning, or additional supervisory signals, greatly increasing model complexity. In contrast, PolyRad employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale. It bypasses the need for complex multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
Figure 2: Overview of PolyRad. PolyRad transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features. Learnable prototypes extract patterns in the configurations of radar point clouds and are used to compute aggregated radar features. The geometry-aware MDE features are fused with the radar features to yield a unified scene representation, which is used to predict coefficients for polynomial fitting, enabling non-uniform corrections that improve accuracy beyond simple scale-and-shift transformations.
3 Method Formulation
We aim to reconstruct a 3D scene from an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a synchronized radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$ by fitting a scaleless depth map from a pretrained MDE foundation model to a metric depth map. Our method leverages MDE to seed scaleless depth predictions and estimates metric depth through polynomial fitting guided by cross-modal features that encode spatial correspondences between the 3D radar points and the dense scaleless depth predictions.
As MDE models are trained on millions of images, they are robust to typical visual nuisance variability, spanning from illumination changes and object appearances to viewpoint shifts. To circumvent the need to collect many large-scale synchronized radar-camera datasets, we begin with a frozen pretrained MDE model $M$, which serves as a learned geometric prior, to infer a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the image. We then encode the point cloud $C$ and the scaleless depth map $z$ separately. Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$:
$\hat{d}\,=\,\hat{c}_{0}+\sum\nolimits_{i=1}^{N}\hat{c}_{i}z^{i}.$ (1)
3.1 Motivation
Existing methods [60, 51, 58, 18, 8, 59] refine MDE predictions by applying simple scale-and-shift transformations. However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths. A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other. Such cross-object misalignments cannot be corrected by a single global scaling factor (see Fig. 1). This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map. Lower-order coefficients (including scale and shift) capture the global scene layout, while higher-order coefficients focus on local depth adjustments, correcting cross-object misalignments and fine-grained MDE errors.
An intuitive understanding of polynomial fitting can be drawn from the geometric perspective of depth transformations. A simple scale and shift operation, mathematically represented as $\hat{a}z+\hat{b}$, uniformly stretches or compresses the entire reconstruction by the same factor across all depth levels. While effective for coarse global corrections, such a linear transformation has zero inflection points, limiting its flexibility to address cross-region misalignments in relative depth predictions. In contrast, the total degree of our polynomial fitting method determines the maximum number of potential inflection points, where the curvature of the depth transformation can change. For a polynomial $f(z)=\sum_{i=0}^{N}\hat{c}_{i}\,z^{i}$, an inflection point $z^{*}$ occurs if
[…]
indicating where the second derivative changes sign. Our method learns inflection points to model transition regions where the curvature of the MDE error changes, allowing necessary non-uniform corrections.
nuScenes ZJU View-of-Delft
       Distance Method MAE RMSE MAE RMSE MAE RMSE
       50m Lin [26] [IROS ’20] 1598.2 3790.1 1576.3 3464.7 2894.9 5507.7
RadarNet [45] [CVPR ’23] 1727.7 3746.8 1430.5 3250.8 2944.8 6113.2
SparseBeatsDense [24] [ECCV ’24] 1524.5 3567.3 1424.4 3267.5 2909.9 5746.4
GET-UP [49] [WACV ’25] 1241.0 2857.0 1483.9 3220.5 2330.0 4565.0
RadarCam-Depth [25] [ICRA ’24] 1286.1 2964.3 1067.5 2817.4 1895.1 4458.7
PolyRad (Ours) 1146.7 2695.6 578.0 1108.6 1293.7 2420.6
       70m Lin [26] [IROS ’20] 1897.8 4558.7 1701.7 3853.3 3312.3 6480.5
RadarNet [45] [CVPR ’23] 2073.2 4590.7 1543.8 3655.3 3428.9 7331.9
SparseBeatsDense [24] [ECCV ’24] 1822.9 4303.6 1520.0 3593.4 3408.3 6914.7
GET-UP [49] [WACV ’25] 1541.0 3657.0 1651.5 3711.7 2758.0 5678.0
RadarCam-Depth [25] [ICRA ’24] 1587.9 3662.5 1157.0 3117.7 2095.4 4944.6
PolyRad (Ours) 1402.9 3269.7 603.7 1154.9 1442.9 2803.8
       80m Lin [26] [IROS ’20] 1988.4 4841.1 1739.5 3993.9 3468.1 6910.3
RadarNet [45] [CVPR ’23] 2179.3 4898.7 1578.4 3804.2 3597.0 7809.2
SparseBeatsDense [24] [ECCV ’24] 1927.0 4609.6 1548.4 3708.1 3584.0 7375.2
GET-UP [49] [WACV ’25] 1632.0 3932.0 1699.7 3882.6 2917.3 6145.1
RadarCam-Depth [25] [ICRA ’24] 1689.7 3948.0 1183.5 3229.0 2227.4 5385.8
PolyRad (Ours) 1454.1 3488.9 629.6 1171.3 1500.1 3951.8
Table 1: Quantitative results on nuScenes, ZJU-4DRadarCam, and View-of-Delft test sets evaluated with various maximum depths.
3.2 Radar Processing
We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where $N_{C}$ denotes the number of points and each point is represented using its coordinates in three-dimensional camera space. Radar point clouds, while noisy and sparse, provide metric depth measurements. We encode $C$ using a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these prototypes can be interpreted as capturing diverse spatial and geometric properties present in the radar point cloud. Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes. Unlike existing works [45, 25] that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud. This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
The next step is to match the learnable prototypes to the most relevant radar features for a given input. We allow each prototype to dynamically attend to the radar points that provide the most relevant spatial and geometric signals, thereby aggregating a sparse point cloud into a global scene-descriptive representation. To achieve this, we project the point-wise radar features $F_{r}$ into a key-value space, where they can be efficiently retrieved by the prototypes $P$. We compute the aggregated radar features $R$:
$R=\text{softmax}\left(\frac{P\times(\psi_{r}^{K}(F_{r}))^{T}}{\sqrt{c_{r}}}\right)\psi_{r}^{V}(F_{r}),$ (3)
where $\psi_{r}^{K}$ and $\psi_{r}^{V}$ are MLPs that project the point-wise radar features $F_{r}$. Comparatively, applying depth completion methods that densify a sparse projection of points using surrounding context [31, 36] to a radar point cloud results in poor performance (see Supp. Mat). This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution [45].
Next, the scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded using a depth encoder, $f_{z}$, to produce depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$. We capitalize on the invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—by extracting features from the MDE depth maps, which leads to features encoding shapes and surfaces. We note that variations in the shape of objects or their geometry are comparatively smaller than variations in visual appearance, which are highly sensitive to transient factors such as perspective and lighting. We then learn spatial correspondences between the depth features $Z$ and the radar features $R$ to produce a unified scene representation that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features. An important insight is that our depth encoder $f_{z}$ yields more stable geometric context by matching radar point configurations to shapes and surfaces that remain relatively consistent across scenes, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
To this end, we fuse the depth features $Z$ and the aggregated radar features $R$ to yield a unified scene representation
$S=\text{softmax}\left(\frac{Z\times(\psi_{R}^{K}(R))^{T}}{\sqrt{c_{r}}}\right)\psi_{R}^{V}(R),$ (4)
where $\psi_{R}^{K}$ and $\psi_{R}^{V}$ are MLPs that project the aggregated radar features $R$. The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield
$\bar{S}=\text{GAP}(f_{s}(S)),$ (5)
which is a $c_{s}$-dimensional feature vector. The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context. This final scene representation $\bar{S}$ is then fed into an MLP $\psi_{s}$, which predicts the $N+1$ polynomial coefficients as a vector $\hat{c}$:
$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (6)
These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities. Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and inter-region corrections, resulting in a high-fidelity metric-scale depth map.
Figure 3: Qualitative comparison on nuScenes. Baseline methods GET-UP and RadarCam-Depth (RC-D) fail to correctly reconstruct the 3D scene, resulting in objects with incorrect or missing depth predictions. While raw MDE outputs reasonable scaleless reconstructions, several cross-object misalignments are apparent. In contrast, PolyRad leverages polynomial fitting to infer high-fidelity depth predictions.
3.3 Polynomial Fitting
Given a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ predicted by the MDE foundation model, our goal is to transform it into a metric depth map $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ using the polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\dots,\hat{c}_{N}\}$ predicted by our network (Sec. 3.2).
Formally, we express the final depth map $\hat{d}$ as:
$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W\end{split}$ (7)
This polynomial formulation applies the learned coefficients to progressively higher powers of the scaleless depth map, enabling complex, non-linear transformations that go beyond simple scale-and-shift adjustments. Each pixel’s final depth is thus computed by combining multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
This is in contrast to directly mapping to a metric-scale depth map [45, 25], where the degrees of freedom comprise every predicted pixel. Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system. As the radar point cloud (hundreds of points) only occupies a small subset of the image space, the solution is underdetermined and often irregular (see Fig. 3). Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab. 3). PolyRad offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
In terms of overhead, compared to predicting scale and shift coefficients for a linear transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost: an MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Supp. Mat).
Geometric Intuition. Applying polynomial fitting to the scaleless depth map can be viewed as a flexible, depth-dependent correction mechanism. In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates. Lower-order terms ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) capture broad scene structure, ensuring that near and far objects maintain reasonable depth relationships. Higher-order terms ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) refine depth predictions by introducing localized adjustments that correct subtle variations, such as mid-range depth discrepancies or fine object boundaries.
The concept of inflection points provides an intuitive understanding of this approach. Each additional polynomial order introduces more potential inflection points, allowing the depth transformation to shift curvature where needed. This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction. Higher-order terms magnify small discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct. In this way, the degree of the polynomial, selected as a hyperparameter, directly reflects the capacity of our model to perform non-uniform depth-dependent adjustments.
The sign of the predicted polynomial coefficients furthers our interpretation. Positive coefficients for higher-order terms push depth values outward while negative ones pull them inward, effectively expanding or contracting selected depth intervals to correct local errors. In doing so, these coefficients shape the curvature of the polynomial, dictating where and how inflection points arise. For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations. By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (see Fig. 4) for how our method adjusts depth predictions at different scales.
Figure 4: PolyRad leverages higher-degree polynomials to reduce error by correcting depth misalignments using inflection points.
3.4 Loss Function
We employ a loss function comprising three terms to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
$\mathcal{L}=\lambda_{1}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{2}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{3}\left\|1-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (8)
where $d_{\text{gt}}\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
The first two terms, $L_{1}$ and $L_{2}$ losses, ensure that our predicted depth map $\hat{d}$ closely matches the ground truth $d_{\text{gt}}$ by penalizing discrepancies. The $L_{1}$ term is less sensitive to outliers and thus promotes robustness, while the $L_{2}$ term penalizes larger errors more significantly.
The novel component of our loss function lies in the third term, which encourages the first derivative of the predicted depth map $\hat{d}$ with respect to the input scaleless depth $z$ to generally be close to 1. Formally, we define:
$\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}(x,y)=\sum_{i=1}^{N}\,i\,\hat{c}_{i}\,z(x,y)^{i-1}.$ (9)
This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression [1]. Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth. The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
However, polynomial fitting introduces a large function space, with expressiveness increasing as the polynomial degree $N$ grows, potentially disrupting local monotonic depth ordering if not properly constrained. Our regularization term addresses this by encouraging that, on average across the scene, the slope of our polynomial transformation remains close to 1, thus preserving monotonicity within local regions. At the same time, since there are fewer measurements between objects, PolyRad is free to correct cross-object misalignments while preserving local depth ordinality.
A first derivative close to 1 means that for every incremental change in the input scaleless depth $z$, there is an approximately equal incremental change in the predicted metric depth $\hat{d}$. This constraint ensures that our model does not overfit to noisy radar data or introduce spurious corrections by avoiding highly non-linear transformations that could disrupt the monotonicity of depth within local regions. Furthermore, this regularization term mitigates oscillatory polynomial curves, a known challenge when using higher-degree polynomials [3]. Without this constraint of regularity, the polynomial may excessively oscillate due to overfitting, leading to erratic depth predictions in some regions.
4 Experiments
Baseline Methods. We consider five recent radar-camera depth estimation baselines: RadarNet [45], Lin [26], SparseBeatsDense [24], GET-UP [49], and RadarCam-Depth [25]. Notably, RadarCam-Depth also leverages radar points to rescale initial scaleless MDE predictions.
Datasets. We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Supp. Mat for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
4.1 Main Results
PolyRad reduces MAE by 12.1% and RMSE by 10.5% on nuScenes, 46.8% and 62.4% on ZJU, and 31.8% and 38.5% on VoD, achieving SOTA results for all datasets (see Tab. 1).
We qualitatively compare PolyRad against all baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (see Fig. 3). PolyRad more accurately predicts metric scale and exhibits fewer global misalignments. Furthermore, it is able to extrapolate depth predictions of object surfaces from learned photometric priors, an ability inherited from the backbone MDE model (UniDepth) trained on large-scale datasets. Our polynomial fitting approach then non-uniformly corrects depth discrepancies across the scene, leading to significant performance improvement over initial MDE predictions. Introducing multiple degrees of freedom enables PolyRad to correct misalignments of objects relative to each other, improving accuracy by correcting consistent over- or under-estimations in local regions. For instance, the raw MDE prediction for Image C incorrectly places the bus stop roof at around the same depth as the more distant tree branch, as highlighted on the right. Additionally, MDE does not accurately infer the boundary between the curb and asphalt, again placing both at erroneously similar depths. With a learned polynomial transformation, PolyRad corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure. In Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths. Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts. This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the construction crane arms in Image B. GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
nuScenes
Method MAE RMSE
DPT [41] 5188.2 6884.5
Depth Anything [56] 2404.9 4851.1
Depth Pro [4] 3835.0 6600.3
UniDepth [37] 2129.8 4887.7
RadarCam-Depth w/ DPT 1689.7 3948.0
RadarCam-Depth w/ Depth Anything 1953.6 5107.8
RadarCam-Depth w/ Depth Pro 3417.1 6462.0
RadarCam-Depth w/ UniDepth 1872.0 4321.2
PolyRad w/ DPT 1525.6 3745.0
PolyRad w/ Depth Anything 1515.1 3719.4
PolyRad w/ Depth Pro 1627.7 4143.9
PolyRad w/ UniDepth 1454.1 3488.9
Table 2: MDE backbone comparative studies on nuScenes.
4.2 Comparative Studies
MDE Models. We evaluate PolyRad using four monocular depth estimation (MDE) backbones: DPT [41], Depth Anything [56], Depth Pro [4], and UniDepth [37]. DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see Tab. 2 and Supp. Mat), and thus we process them the same way as scaleless depth.
Tab. 2 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and PolyRad. Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements. In contrast, PolyRad consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting. For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results. Among all configurations, PolyRad w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 30.2% and over RadarCam-Depth w/ UniDepth by 20.8%.
Polynomial Degree. As polynomial degree—selected as a hyperparameter—increases, so does the number of inflection points, allowing PolyRad to correct misalignments between the many scene elements in Fig. 4, including cars, trucks, trees, buildings, and traffic lights. This is reflected in the successively improved error maps. Tab. 3 quantifies the performance gains achieved with higher-degree polynomials. At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
Ablations. We evaluate the impact of key components of PolyRad in Tab. 4. Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations. Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features. Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
nuScenes ZJU
Polynomial Degree MAE RMSE MAE RMSE
1 (Scale + Shift) 2156.8 4491.3 1078.2 2405.5
2 1715.2 3840.6 901.0 1822.7
4 1482.7 3510.1 791.2 1516.4
6 1466.9 3496.0 670.3 1297.9
8 1454.1 3488.9 629.6 1171.3
10 1463.7 3494.5 643.3 1184.5
Table 3: Sensitivity study of polynomial degree hyperparameter.
nuScenes VoD
Ablated Component MAE RMSE MAE RMSE
learnable prototypes 1615.5 3629.0 1619.3 4162.7
cross-modality att. 2238.8 4817.4 2147.9 4926.1
monotonicity loss 1921.1 4399.5 1924.5 4660.3
no ablations 1454.1 3488.9 1500.1 3951.8
Table 4: Ablation studies of architecture and loss components.
5 Discussion
Limitations. PolyRad requires tuning the polynomial degree as a hyperparameter, which may vary across datasets. Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
Summary. We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth. Our principled and efficient approach demonstrates that a fundamental insight—shifting from simple scale-and-shift to flexible, radar-guided polynomial transformations—can outperform more complex systems.

### v2

@@
 Abstract
-We propose PolyRad, a novel radar-guided depth estimation method that introduces polynomial fitting to transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps.
-Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a simple yet fundamental insight: using polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges.
-Although MDE models often infer reasonably accurate local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale-and-shift transformation insufficient given three or more of these regions.
-In contrast, PolyRad generalizes beyond linear transformations and is able to correct such misalignments by introducing inflection points.
-Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces monotonicity via first-derivative regularization.
-PolyRad achieves state-of-the-art performance on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets, outperforming existing methods by 30.3% in MAE and 37.2% in RMSE.
+We propose POLAR, a novel radar-guided depth estimation method that introduces polynomial fitting to efficiently transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps.
+Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a fundamental insight: although MDE models often infer reasonable local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale and shift (affine) transformation insufficient given three or more of these regions.
+To address this limitation, we use polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges.
+In this way, POLAR generalizes beyond affine transformations and is able to correct such misalignments by introducing inflection points.
+Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces local monotonicity via first-derivative regularization.
+POLAR achieves state-of-the-art performance across three datasets, outperforming existing methods by an average of 24.9% in MAE and 33.2% in RMSE, while also achieving state-of-the-art efficiency in terms of latency and computational cost.
 1 Introduction
-Metric 3D reconstruction is critical for spatial tasks such as self-driving [34, 14], where it is necessary for one to perceive the structure of the 3D environment in order to navigate.
-In many such systems, multiple sensors—including cameras, lidar, and radar provide complementary information about the 3D scene.
-While lidar sensors offer dense, accurate point clouds, they are expensive and not widely available [39].
-In contrast, radar sensors, particularly millimeter wave (mmWave) radars [20], return only about a hundred points per frame and are noisier [15].
-Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles [10].
-Figure 1: Once an MDE model predicts incorrect relative depths between three or more objects, a linear scale-and-shift (dashed) cannot resolve this misalignment.
-Instead, we learn a polynomial fit with inflection points, enabling flexible non-uniform corrections.
+Metric 3D reconstruction is critical for spatial tasks such as self-driving (Maier et al., 2012; Gupta et al., 2021), where it is necessary for one to perceive the structure of the 3D environment in order to navigate.
+In many such systems, multiple sensors—including cameras, lidar, and radar—provide complementary information about the 3D scene.
+While lidar sensors offer dense and precise point clouds, they are expensive and not widely available (Raj et al., 2020).
+In contrast, radar sensors, particularly millimeter wave (mmWave) radars (Iizuka et al., 2003), return only about a hundred points per frame and are noisier (Han et al., 2023).
+Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles (Eichelberger & McCartt, 2016).
 Unless camera baselines are known, images offer scaleless reconstruction.
 However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint.
-Training robust models that can generalize across these factors often demands large-scale datasets, which are costly and time-intensive to collect.
-As a result, leveraging pretrained monocular depth estimation (MDE) models [41, 56, 4, 37] has emerged as a practical alternative.
-However, as these models only take in a single image, where reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or metric depth that lacks the fidelity needed (see Tab.
-2) for applications requiring accurate metric-scale reconstruction, e.g., navigation.
-Existing approaches [60, 51, 58, 18, 8, 59] attempt to transform these MDE predictions into accurate metric depth maps using a global linear (scale-and-shift) transformation.
+Training robust models that can generalize across these factors often demands large-scale datasets, which are costly to collect.
+As a result, leveraging pretrained monocular depth estimation (MDE) foundation models (Ranftl et al., 2021; Yang et al., 2024; Bochkovskii et al., 2024; Piccinelli et al., 2025) emerges as a practical alternative.
+However, as monocular 3D reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or depth that lacks the fidelity needed for applications demanding accurate metric-scale reconstruction, such as mapping and navigation.
+Existing approaches (Yin et al., 2023; Viola et al., 2024; Hu et al., 2024; Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025) attempt to transform these MDE predictions into accurate metric depth maps using a global affine (scale-and-shift) transformation.
 This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene.
 While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other.
-Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (see Fig.
+Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (Fig.
 1).
-Empirically, a learned linear fit on a challenging dataset such as nuScenes performs worse than even a deterministic estimator like median scaling (see Tab.
-2), indicating susceptibility to outliers.
-In this work, we challenge the common misconception that scale ambiguity in MDE is only up to an unknown global scale and shift.
-Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, thereby enabling more flexible non-uniform adjustments.
-Our approach, PolyRad, transforms the predictions of pretrained MDE models by: learning prototypical patterns in the configuration of radar points, establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and predicting polynomial coefficients that adaptively fit scaleless depth into metric depth.
-This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and inter-region misalignments.
-While one may adjust every prediction (pixel) of an MDE model [25], the number of degrees of freedom is the total number of pixels.
+Figure 1: If an MDE model predicts incorrect relative depths between three or more objects, an affine scale-and-shift (dashed) cannot resolve this misalignment.
+POLAR (solid) overcomes this limitation by learning an NN’th-order polynomial fit with up to N2N-2 inflection points.
+In this work, we challenge the common assumption that scale ambiguity in MDE is only up to an unknown global scale and shift.
+Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, enabling more flexible non-uniform adjustments.
+Our approach, POLAR, transforms the predictions of pretrained MDE models by (i) learning prototypical patterns in the configuration of radar points, (ii) establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and (iii) predicting polynomial coefficients that adaptively fit scaleless depth into metric depth.
+This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and cross-region misalignments.
+While one may adjust every prediction (pixel) of an MDE model (Li et al., 2024a), the number of degrees of freedom is the total number of pixels.
 Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem.
 On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular.
-To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models, making our approach a novel solution not only for radar-guided depth estimation but also for any task that leverages foundation models and refines their predictions using additional inputs as guidance.
+To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models.
 However, learning such a flexible transformation is challenging because higher-order polynomials introduce a vast function space with many free parameters.
@@
 This expressiveness can inadvertently lead to harmful non-monotonic transformations—where the relative depth orderings of points are incorrectly reversed—if not properly constrained.
-We address this by introducing a regularization term that enforces the first derivative of the predicted metric depth with respect to the input MDE depth to generally remain close to 1, ensuring that incremental changes in the input MDE depth result in approximately equal changes in the predicted depth within local regions where MDE reconstructs structure with high fidelity.
-In essence, this regularizer encourages a smooth, monotonic transformation, mitigating unstable oscillations that can arise from overfitting high-degree polynomials while still enabling necessary corrections for cross-region misalignments.
-Our contributions: We propose (1) a novel radar-guided depth estimation method that leverages polynomial fitting to transform scaleless MDE predictions into accurate metric depth maps.
-Our approach (2) introduces a fundamental insight: using polynomial coefficients predicted from radar points to enable non-uniform, depth-dependent corrections.
-We present (3) a principled geometric approach, where polynomial transformations naturally introduce inflection points that can correct misalignments between local areas that a global scale-and-shift transformation cannot.
-We design (4) a novel training objective that encourages monotonicity through first-derivative as regularization, ensuring structural consistency while allowing inter-region depth adjustments.
-Finally, (5) extensive experiments on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets demonstrate that PolyRad achieves the state of the art, outperforming existing methods by 11.3%, 54.6%, and 35.2%, respectively.
+We address this by introducing a regularization term that enforces the predicted metric depth with respect to the input MDE depth to remain approximately monotonically increasing.
+This ensures that incremental changes in the input depth result in proportional changes in the predicted depth within local regions where MDE reconstructs structure up to a relative scale.
+In essence, this regularizer encourages a piecewise monotonic transformation, mitigating unstable oscillations that can arise from overfitting high-degree polynomials while still enabling necessary corrections for cross-region misalignments.
+Our contributions: We propose (1) POLAR, a novel POLynomial fitting method that leverages complementary radAR guidance to transform scaleless monocular depth into accurate metric depth.
+Our approach (2) introduces a fundamental insight: using polynomial coefficients predicted from a learned multimodal representation to enable non-uniform corrections.
+We present (3) a principled geometric formulation, where polynomial transformations introduce inflection points that can correct misalignments between local areas that an affine transformation cannot.
+We design (4) a novel training objective that encourages monotonicity through first-derivative as regularization, preserving local ordinality while allowing cross-region adjustments.
+Finally, (5) extensive experiments demonstrate that POLAR achieves the state of the art in both performance and efficiency, outperforming existing methods by an average of 29.1% and simultaneously delivering real-time processing of over 40 fps.
 2 Related Work
-2.1 Monocular Depth Estimation
-With recent advances in the scalability of neural networks [9, 5], monocular depth estimation (MDE) models can infer relative depth from a single image across diverse, unseen domains [22, 40, 55, 56, 37, 4], benefiting from large-scale datasets.
-Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation [16, 2], uncertainty estimation [38], optical flow [62], and visual odometry [46], to improve relative depth estimation.
+Monocular Depth Estimation.
+With recent advances in the scalability of neural networks (Dosovitskiy et al., 2021; Caron et al., 2021), monocular depth estimation (MDE) models can infer scaleless relative depth from a single image across diverse, unseen domains (Ranftl et al., 2020; Yang et al., 2024; Piccinelli et al., 2025; Bochkovskii et al., 2024), benefiting from large-scale datasets.
+Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation (Hoyer et al., 2023; Bian et al., 2019), uncertainty estimation (Poggi et al., 2020), optical flow (Zhao et al., 2020), and visual odometry (Song et al., 2023), to improve relative depth estimation.
 However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem.
-Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity [51].
-One way to address this limitation is to incorporate range-sensing modalities such as lidar [21, 11], radar [45], or visual-inertial odometry [52].
-2.2 Range-Camera Depth Estimation
-Metric-scale depth estimation often requires recovering a dense depth map from sparse range measurements, guided by complementary camera inputs.
-Image-guided depth completion, such as lidar-camera depth estimation, leverages the strengths of each modality: images offer dense visual context and structural priors, while sparse depth provides absolute metric scale to resolve depth ambiguity.
-Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers [32, 33], late fusion, which maintains independent branches for different sensors before merging outputs [57, 54, 50, 12, 43], and multi-scale fusion, which captures both local details and global scene structure [23].
-U-Net-like architectures with CSPN-based refinement [6] have been widely used for coarse-to-fine depth completion [17, 27, 36], with further improvements from adaptive convolutions [7], deformable convolutions [36, 53], and transformer-based attention mechanisms [42, 61].
-2.3 Radar-Camera Depth Estimation
-While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions [39].
-In contrast, radar provides a cost-effective alternative [19], offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle [35, 47, 44].
-The ubiquity of radar sensors [10] in existing automotive and robotic platforms further supports their integration into depth estimation pipelines.
+Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity (Viola et al., 2024).
+One way to address this limitation is to incorporate range-sensing modalities such as lidar (Jaritz et al., 2018; Ezhov et al., 2024), radar (Singh et al., 2023), or visual-inertial odometry (Wong et al., 2020).
+Image-Guided Depth Completion.
+Most often studied in the context of lidar-camera depth estimation, image-guided depth completion leverages the strengths of each modality: images offer dense visual context and structural priors, while lidar points provide absolute metric scale to resolve depth ambiguity.
+Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers (Ma & Karaman, 2018; Ma et al., 2018), late fusion, where inputs are processed by independent branches (Yan et al., 2021; Rim et al., 2025), and multi-scale fusion, which captures both local details and global scene structure (Li et al., 2020).
+U-Net-like architectures have been widely used for coarse-to-fine depth completion (Hu et al., 2021; Lin et al., 2022), with improvements from deformable convolutions (Park et al., 2020; Xu et al., 2020), and attention mechanisms (Rho et al., 2022; Zhang et al., 2023).
+Radar-Camera Depth Estimation.
+While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions (Raj et al., 2020).
+In contrast, radar provides a cost-effective alternative (Hunt et al., 2024), offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle (Paek et al., 2022; Srivastav & Mandal, 2023).
+The ubiquity of radar sensors (Eichelberger & McCartt, 2016) in existing automotive and robotic platforms further supports their integration into depth estimation pipelines.
 Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
-Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity [45] of radar point clouds.
-[26] uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map.
-[29] leverages Doppler velocity and optical flow to associate radar points with image pixels.
-Other methods refine radar depth using neighborhood pixel associations [30], height-extension of radar points to address elevation ambiguity [28], or by extracting self-supervisory signals from the radar point cloud [13].
-[45] introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion.
-[24] mitigates distribution artifacts using sparse supervision, while [48] employs a two-stage confidence-driven approach.
-[49] uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features.
-Like our method, [25] starts with an initial MDE prediction but refines it through quasi-dense scale estimation, followed by local scale adjustment, rather than through a learned scene-fitting approach.
-While these existing methods employ complex architectures to refine radar-derived spatial information, they all operate within the paradigm of completion or confidence-driven fusion.
-They often rely on multi-stage training, explicit radar-pixel association learning, or additional supervisory signals, greatly increasing model complexity.
-In contrast, PolyRad employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale.
-It bypasses the need for complex multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
-Figure 2: Overview of PolyRad.
-PolyRad transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features.
-Learnable prototypes extract patterns in the configurations of radar point clouds and are used to compute aggregated radar features.
-The geometry-aware MDE features are fused with the radar features to yield a unified scene representation, which is used to predict coefficients for polynomial fitting, enabling non-uniform corrections that improve accuracy beyond simple scale-and-shift transformations.
+Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity (Singh et al., 2023) of radar point clouds.
+(Lin et al., 2020) uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map.
+(Long et al., 2021) leverages Doppler velocity and optical flow to associate radar points with image pixels.
+(Lo & Vandewalle, 2021) refines radar depth using height-extension of radar points to address elevation ambiguity.
+(Singh et al., 2023) introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion.
+(Li et al., 2024b) mitigates distribution artifacts using sparse supervision, while (Sun et al., 2024) employs a two-stage confidence-driven approach.
+GET-UP (Sun et al., 2025) uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features.
+Like our method, RadarCam-Depth (Li et al., 2024a) and TacoDepth (Wang et al., 2025) also begin with an initial MDE prediction, but refine it and decode dense depth directly from fused features, rather than employing a learned scene-fitting approach.
+Existing methods operate within the paradigm of completion or direct decoding of fused features, and they often rely on multi-stage training and explicit radar-pixel association learning, increasing model complexity and computational overhead.
+In contrast, POLAR employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale.
+It bypasses the need for multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
 3 Method Formulation
@@
 We then encode the point cloud $C$ and the scaleless depth map $z$ separately.
-Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$:
-$\hat{d}\,=\,\hat{c}_{0}+\sum\nolimits_{i=1}^{N}\hat{c}_{i}z^{i}.$ (1)
+Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$.
+Figure 2: Method Overview.
+POLAR transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features.
+Learnable prototypes extract patterns in the configurations of radar point clouds and are used to aggregate spatially-informed radar features.
+The geometry-aware MDE features are fused with the radar features via a learnable soft-correspondence module to yield a unified scene representation that is used to predict polynomial coefficients for fitting.
+This enables non-uniform corrections that improve accuracy beyond affine transformations.
 3.1 Motivation
-Existing methods [60, 51, 58, 18, 8, 59] refine MDE predictions by applying simple scale-and-shift transformations.
+Existing methods (Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025; Viola et al., 2024; Hu et al., 2024; Yin et al., 2023) refine MDE predictions by applying simple scale-and-shift transformations.
 However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths.
 A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other.
-Such cross-object misalignments cannot be corrected by a single global scaling factor (see Fig.
+Such cross-region misalignments cannot be corrected by a single global scaling factor (see Fig.
 1).
 This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
-To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map.
+To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map via summation.
 Lower-order coefficients (including scale and shift) capture the global scene layout, while higher-order coefficients focus on local depth adjustments, correcting cross-object misalignments and fine-grained MDE errors.
 An intuitive understanding of polynomial fitting can be drawn from the geometric perspective of depth transformations.
-A simple scale and shift operation, mathematically represented as $\hat{a}z+\hat{b}$, uniformly stretches or compresses the entire reconstruction by the same factor across all depth levels.
-While effective for coarse global corrections, such a linear transformation has zero inflection points, limiting its flexibility to address cross-region misalignments in relative depth predictions.
+An affine (scale-and-shift) operation, mathematically represented as $\hat{a}z+\hat{b}$, uniformly stretches or compresses the entire reconstruction by the same factor across all depth levels.
+While effective for coarse global corrections, such an affine transformation has zero inflection points, limiting its flexibility to address cross-region misalignments in relative depth predictions.
 In contrast, the total degree of our polynomial fitting method determines the maximum number of potential inflection points, where the curvature of the depth transformation can change.
 For a polynomial $f(z)=\sum_{i=0}^{N}\hat{c}_{i}\,z^{i}$, an inflection point $z^{*}$ occurs if
-[…]
-indicating where the second derivative changes sign.
-Our method learns inflection points to model transition regions where the curvature of the MDE error changes, allowing necessary non-uniform corrections.
-nuScenes ZJU View-of-Delft
-       Distance Method MAE RMSE MAE RMSE MAE RMSE
-       50m Lin [26] [IROS ’20] 1598.2 3790.1 1576.3 3464.7 2894.9 5507.7
-RadarNet [45] [CVPR ’23] 1727.7 3746.8 1430.5 3250.8 2944.8 6113.2
-SparseBeatsDense [24] [ECCV ’24] 1524.5 3567.3 1424.4 3267.5 2909.9 5746.4
-GET-UP [49] [WACV ’25] 1241.0 2857.0 1483.9 3220.5 2330.0 4565.0
-RadarCam-Depth [25] [ICRA ’24] 1286.1 2964.3 1067.5 2817.4 1895.1 4458.7
-PolyRad (Ours) 1146.7 2695.6 578.0 1108.6 1293.7 2420.6
-       70m Lin [26] [IROS ’20] 1897.8 4558.7 1701.7 3853.3 3312.3 6480.5
-RadarNet [45] [CVPR ’23] 2073.2 4590.7 1543.8 3655.3 3428.9 7331.9
-SparseBeatsDense [24] [ECCV ’24] 1822.9 4303.6 1520.0 3593.4 3408.3 6914.7
-GET-UP [49] [WACV ’25] 1541.0 3657.0 1651.5 3711.7 2758.0 5678.0
-RadarCam-Depth [25] [ICRA ’24] 1587.9 3662.5 1157.0 3117.7 2095.4 4944.6
-PolyRad (Ours) 1402.9 3269.7 603.7 1154.9 1442.9 2803.8
-       80m Lin [26] [IROS ’20] 1988.4 4841.1 1739.5 3993.9 3468.1 6910.3
-RadarNet [45] [CVPR ’23] 2179.3 4898.7 1578.4 3804.2 3597.0 7809.2
-SparseBeatsDense [24] [ECCV ’24] 1927.0 4609.6 1548.4 3708.1 3584.0 7375.2
-GET-UP [49] [WACV ’25] 1632.0 3932.0 1699.7 3882.6 2917.3 6145.1
-RadarCam-Depth [25] [ICRA ’24] 1689.7 3948.0 1183.5 3229.0 2227.4 5385.8
-PolyRad (Ours) 1454.1 3488.9 629.6 1171.3 1500.1 3951.8
-Table 1: Quantitative results on nuScenes, ZJU-4DRadarCam, and View-of-Delft test sets evaluated with various maximum depths.
-3.2 Radar Processing
-We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where $N_{C}$ denotes the number of points and each point is represented using its coordinates in three-dimensional camera space.
+$f^{\prime\prime}(z^{*})\,=\,\sum\nolimits_{i=2}^{N}i(i-1)\,\hat{c}_{i}\,(z^{*})^{i-2}\,=\,0,$ (1)
+indicating where the second derivative changes sign. By learning inflection points, we can model transition regions in the curvature of the MDE error, allowing necessary non-uniform corrections.
+3.2 Representation Learning and Fusion
+Radar Processing.
+We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where each of the $N_{C}$ points is represented by its $(x,y,z)$ coordinates in three-dimensional camera space.
 Radar point clouds, while noisy and sparse, provide metric depth measurements.
-We encode $C$ using a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
-To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these prototypes can be interpreted as capturing diverse spatial and geometric properties present in the radar point cloud.
+We concatenate a sinusoidal 3D positional embedding $\phi_{3D}(x,y,z)$ to $C$, and feed the resulting representation into a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
+Radar Aggregation.
+To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these $N_{P}$ prototypes capture diverse spatial and geometric properties present in the radar point cloud.
 Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes.
-Unlike existing works [45, 25] that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud.
+Unlike existing works (Singh et al., 2023; Li et al., 2024a) that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud.
 This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
-The next step is to match the learnable prototypes to the most relevant radar features for a given input.
-We allow each prototype to dynamically attend to the radar points that provide the most relevant spatial and geometric signals, thereby aggregating a sparse point cloud into a global scene-descriptive representation.
-To achieve this, we project the point-wise radar features $F_{r}$ into a key-value space, where they can be efficiently retrieved by the prototypes $P$.
-We compute the aggregated radar features $R$:
-$R=\text{softmax}\left(\frac{P\times(\psi_{r}^{K}(F_{r}))^{T}}{\sqrt{c_{r}}}\right)\psi_{r}^{V}(F_{r}),$ (3)
-where $\psi_{r}^{K}$ and $\psi_{r}^{V}$ are MLPs that project the point-wise radar features $F_{r}$.
-Comparatively, applying depth completion methods that densify a sparse projection of points using surrounding context [31, 36] to a radar point cloud results in poor performance (see Supp.
-Mat).
-This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution [45].
-Next, the scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded using a depth encoder, $f_{z}$, to produce depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$.
-We capitalize on the invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—by extracting features from the MDE depth maps, which leads to features encoding shapes and surfaces.
-We note that variations in the shape of objects or their geometry are comparatively smaller than variations in visual appearance, which are highly sensitive to transient factors such as perspective and lighting.
-We then learn spatial correspondences between the depth features $Z$ and the radar features $R$ to produce a unified scene representation that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features.
-An important insight is that our depth encoder $f_{z}$ yields more stable geometric context by matching radar point configurations to shapes and surfaces that remain relatively consistent across scenes, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
-To this end, we fuse the depth features $Z$ and the aggregated radar features $R$ to yield a unified scene representation
-$S=\text{softmax}\left(\frac{Z\times(\psi_{R}^{K}(R))^{T}}{\sqrt{c_{r}}}\right)\psi_{R}^{V}(R),$ (4)
-where $\psi_{R}^{K}$ and $\psi_{R}^{V}$ are MLPs that project the aggregated radar features $R$.
-The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield
-$\bar{S}=\text{GAP}(f_{s}(S)),$ (5)
-which is a $c_{s}$-dimensional feature vector.
+Next, learnable prototypes are matched to the most relevant radar features for a given input.
+We treat the prototypes as centroids and perform soft clustering over the radar features.
+Each radar point is softly assigned to prototypes according to feature similarity, and the corresponding values are aggregated to yield a global scene-descriptive representation $F_{R}$:
+$D_{ij}=\|P_{j}-\Phi_{r}(F_{r})_{i}\|^{2},\quad F_{R}=\text{softmax}\!\left(-D\,/\,\tau\right)\,\Psi_{r}(F_{r}),$ (2)
+where $\Phi_{r}$, $\Psi_{r}$ are MLPs that project the radar features $F_{r}$, the matrix $D\in\mathbb{R}^{N_{P}\times N_{C}}$ contains the pairwise squared distances between prototypes $\{P_{j}\}_{j=1}^{N_{P}}$ and projected radar features $\{\Phi_{r}(F_{r})_{i}\}_{i=1}^{N_{C}}$, and $\tau$ denotes temperature.
+This clustering-based formulation enables prototypes to capture recurring spatial and geometric patterns in radar point clouds.
+Comparatively, applying lidar depth completion methods, which densify a sparse projection of points using surrounding context, to a radar point cloud results in poor performance (see Sec.
+E).
+This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution (Singh et al., 2023).
+MDE-Radar Fusion.
+The scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded with a learnable depth encoder $f_{z}$, producing depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$.
+These features inherit invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—and therefore primarily encode object-level properties.
+Since variations in the shape of objects and their geometry are generally more stable across scenes than photometric appearance, the resulting depth features provide a more reliable geometric context for fusion with radar features.
+This allows radar point configurations to be matched to consistently observed shapes and surfaces, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
+To this end, we learn soft spatial correspondences between the depth features $Z$ and the radar features $F_{R}$ to construct a unified scene representation $S$ that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features:
+$S=\text{softmax}\left(\frac{(Z+E)\times(\Phi_{R}(F_{R}))^{T}}{\sqrt{c_{r}}}\right)\Psi_{R}(F_{R}),$ (3)
+where $E\in\mathbb{R}^{(H\times W)\times c_{z}}$ is a learned 2D positional embedding, and $\Phi_{R}$, $\Psi_{R}$ are MLPs that project the aggregated radar features $F_{R}$.
+$c_{z}$ is set equal to $c_{r}$ to align the depth and radar features within a common embedding space, facilitating cross-modal fusion, and $S$ is reshaped to be in $\mathbb{R}^{H\times W\times c_{z}}$.
+Predicting Coefficients.
+The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield $\bar{S}=\text{GAP}(f_{s}(S))$, which is a $c_{s}$-dimensional feature vector.
 The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context.
@@
 These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities.
-Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and inter-region corrections, resulting in a high-fidelity metric-scale depth map.
-Figure 3: Qualitative comparison on nuScenes.
-Baseline methods GET-UP and RadarCam-Depth (RC-D) fail to correctly reconstruct the 3D scene, resulting in objects with incorrect or missing depth predictions.
-While raw MDE outputs reasonable scaleless reconstructions, several cross-object misalignments are apparent.
-In contrast, PolyRad leverages polynomial fitting to infer high-fidelity depth predictions.
+Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and cross-region corrections, resulting in a high-fidelity metric-scale depth map.
 3.3 Polynomial Fitting
@@
 Formally, we express the final depth map $\hat{d}$ as:
-$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W\end{split}$ (7)
-This polynomial formulation applies the learned coefficients to progressively higher powers of the scaleless depth map, enabling complex, non-linear transformations that go beyond simple scale-and-shift adjustments.
-Each pixel’s final depth is thus computed by combining multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
-This is in contrast to directly mapping to a metric-scale depth map [45, 25], where the degrees of freedom comprise every predicted pixel.
+$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W.\end{split}$ (5)
+Figure 3: POLAR leverages spatial information from radar points to predict higher-degree polynomial transformations that can correct non-affine errors in MDE predictions.
+This polynomial formulation applies the learned coefficients to successive powers of the scaleless depth map, enabling complex, non-linear transformations that surpass the limitations of affine transformations.
+Each pixel’s final depth is thus computed by summing multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
+This is in contrast to directly mapping to a metric-scale depth map (Singh et al., 2023; Li et al., 2024a), where the degrees of freedom comprise every predicted pixel.
 Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system.
-As the radar point cloud (hundreds of points) only occupies a small subset of the image space, the solution is underdetermined and often irregular (see Fig.
-3).
+As the radar point cloud ($\sim 10^{2}$ points) only occupies a small subset of the image space ($\sim 10^{6}$ pixels), the solution is underdetermined and often irregular (see Fig.
+4).
 Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab.
-3).
-PolyRad offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
-In terms of overhead, compared to predicting scale and shift coefficients for a linear transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost: an MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Supp.
-Mat).
+6).
+POLAR offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
+In terms of overhead, compared to predicting two coefficients (scale and shift) for an affine transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost ($0.01\%$ more FLOPs): the final MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Sec.
+B for the complete derivation).
 Geometric Intuition.
-Applying polynomial fitting to the scaleless depth map can be viewed as a flexible, depth-dependent correction mechanism.
+Applying polynomial fitting to the scaleless depth map is best viewed as a flexible, depth-dependent correction mechanism.
 In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
-By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates.
-Lower-order terms ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) capture broad scene structure, ensuring that near and far objects maintain reasonable depth relationships.
-Higher-order terms ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) refine depth predictions by introducing localized adjustments that correct subtle variations, such as mid-range depth discrepancies or fine object boundaries.
+By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates $z$.
+Lower-order coefficients ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) establish a broad global scale, providing a metric basis for the initially scaleless $z$.
+Higher-order coefficients ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) introduce both low-frequency corrections, e.g., resolving cross-object misalignments (Fig.
+1), as well as high-frequency corrections, e.g., sharpening fine object boundaries (Fig.
+4, Image C).
 The concept of inflection points provides an intuitive understanding of this approach.
@@
 This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction.
-Higher-order terms magnify small discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct.
-In this way, the degree of the polynomial, selected as a hyperparameter, directly reflects the capacity of our model to perform non-uniform depth-dependent adjustments.
+Higher-order terms magnify small (i.e., high-frequency) discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct.
+In this way, the degree of the polynomial, selected as a hyperparameter, governs the capacity of our model to apply non-uniform corrections across the scene.
 The sign of the predicted polynomial coefficients furthers our interpretation.
@@
 For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations.
-By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (see Fig.
-4) for how our method adjusts depth predictions at different scales.
-Figure 4: PolyRad leverages higher-degree polynomials to reduce error by correcting depth misalignments using inflection points.
+By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (e.g., Fig.
+3) for refining scaleless depth into high-fidelity metric depth.
+Table 1: Quantitative results on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft test sets evaluated with various maximum evaluation distances.
+nuScenes ZJU View-of-Delft
+   Distance Method MAE RMSE MAE RMSE MAE RMSE
+   50m RadarNet [CVPR ’23] 1727.7 3746.8 1430.5 3250.8 2944.8 6113.2
+SparseBeatsDense [ECCV ’24] 1524.5 3567.3 1424.4 3267.5 2909.9 5746.4
+GET-UP [WACV ’25] 1241.0 2857.0 1483.9 3220.5 2330.0 4565.0
+RadarCam-Depth [ICRA ’24] 1286.1 2964.3 1067.5 2817.4 1895.1 4458.7
+TacoDepth [CVPR ’25] 1046.8 2487.5 930.2 2477.3 - -
+POLAR (Ours) 1014.4 2475.7 578.0 1108.6 1293.7 2420.6
+   70m RadarNet [CVPR ’23] 2073.2 4590.7 1543.8 3655.3 3428.9 7331.9
+SparseBeatsDense [ECCV ’24] 1822.9 4303.6 1520.0 3593.4 3408.3 6914.7
+GET-UP [WACV ’25] 1541.0 3657.0 1651.5 3711.7 2758.0 5678.0
+RadarCam-Depth [ICRA ’24] 1587.9 3662.5 1157.0 3117.7 2095.4 4944.6
+TacoDepth [CVPR ’25] 1347.1 3152.8 983.1 2779.6 - -
+POLAR (Ours) 1286.1 2947.3 603.7 1154.9 1442.9 2803.8
+   80m RadarNet [CVPR ’23] 2179.3 4898.7 1578.4 3804.2 3597.0 7809.2
+SparseBeatsDense [ECCV ’24] 1927.0 4609.6 1548.4 3708.1 3584.0 7375.2
+GET-UP [WACV ’25] 1632.0 3932.0 1699.7 3882.6 2917.3 6145.1
+RadarCam-Depth [ICRA ’24] 1689.7 3948.0 1183.5 3229.0 2227.4 5385.8
+TacoDepth [CVPR ’25] 1492.4 3324.8 1032.5 2850.3 - -
+POLAR (Ours) 1407.8 3193.5 629.6 1171.3 1500.1 3951.8
 3.4 Loss Function
-We employ a loss function comprising three terms to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
-$\mathcal{L}=\lambda_{1}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{2}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{3}\left\|1-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (8)
+We employ a loss function comprising three terms, each weighted by its $\lambda$, to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
+$\mathcal{L}=\lambda_{L_{1}}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{L_{2}}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{\text{mono}}\left\|\mathbf{1}_{H\times W}-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (6)
 where $d_{\text{gt}}\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
@@
 The $L_{1}$ term is less sensitive to outliers and thus promotes robustness, while the $L_{2}$ term penalizes larger errors more significantly.
-The novel component of our loss function lies in the third term, which encourages the first derivative of the predicted depth map $\hat{d}$ with respect to the input scaleless depth $z$ to generally be close to 1.
-Formally, we define:
+The novel component of our objective lies in the third term, which constrains the first derivative of the predicted depth $\hat{d}$ w.r.t the input scaleless depth $z$ (Eq.
+7) to remain near that of $\hat{d}=z$.
 $\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}(x,y)=\sum_{i=1}^{N}\,i\,\hat{c}_{i}\,z(x,y)^{i-1}.$ (7)
-This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression [1].
+This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression (Barlow et al., 1972).
 Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth.
 The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
-However, polynomial fitting introduces a large function space, with expressiveness increasing as the polynomial degree $N$ grows, potentially disrupting local monotonic depth ordering if not properly constrained.
-Our regularization term addresses this by encouraging that, on average across the scene, the slope of our polynomial transformation remains close to 1, thus preserving monotonicity within local regions.
-At the same time, since there are fewer measurements between objects, PolyRad is free to correct cross-object misalignments while preserving local depth ordinality.
-A first derivative close to 1 means that for every incremental change in the input scaleless depth $z$, there is an approximately equal incremental change in the predicted metric depth $\hat{d}$.
-This constraint ensures that our model does not overfit to noisy radar data or introduce spurious corrections by avoiding highly non-linear transformations that could disrupt the monotonicity of depth within local regions.
-Furthermore, this regularization term mitigates oscillatory polynomial curves, a known challenge when using higher-degree polynomials [3].
-Without this constraint of regularity, the polynomial may excessively oscillate due to overfitting, leading to erratic depth predictions in some regions.
+Polynomial fitting introduces a large function space, with expressiveness growing with degree $N$, which can significantly and detrimentally disrupt local monotonic depth ordering if unconstrained.
+Our regularization term addresses this by preserving local depth ordinality while still allowing corrections of cross-object misalignments.
+This constraint prevents overfitting of nonlinear transformations to noisy radar data, spurious correlations, and outliers, thereby avoiding the oscillatory behavior that is a known challenge with fitting higher-degree polynomials (Bishop, 2006).
+Figure 4: Qualitative results on nuScenes.
+GET-UP and RadarCam-Depth (RC-D) fail to reconstruct entire regions, yielding objects with large depth errors.
+Raw MDE yields reasonable relative reconstructions but suffers from incorrect global scale and cross-object misalignments.
+POLAR leverages polynomial fitting to recover a global scale and correct these misalignments.
+See Fig.
+5 for colorbars.
 4 Experiments
 Baseline Methods.
-We consider five recent radar-camera depth estimation baselines: RadarNet [45], Lin [26], SparseBeatsDense [24], GET-UP [49], and RadarCam-Depth [25].
-Notably, RadarCam-Depth also leverages radar points to rescale initial scaleless MDE predictions.
+We consider five recent radar-camera depth estimation baselines: RadarNet (Singh et al., 2023), SparseBeatsDense (Li et al., 2024b), GET-UP (Sun et al., 2025), RadarCam-Depth (Li et al., 2024a), and TacoDepth (Wang et al., 2025).
+Notably, both RadarCam-Depth and TacoDepth also take scaleless MDE predictions as input, together with radar points, to predict metric depth.
 Datasets.
-We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Supp.
-Mat for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
+We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Sec.
+C for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
 4.1 Main Results
-PolyRad reduces MAE by 12.1% and RMSE by 10.5% on nuScenes, 46.8% and 62.4% on ZJU, and 31.8% and 38.5% on VoD, achieving SOTA results for all datasets (see Tab. 1).
-We qualitatively compare PolyRad against all baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (see Fig.
-3).
-PolyRad more accurately predicts metric scale and exhibits fewer global misalignments.
-Furthermore, it is able to extrapolate depth predictions of object surfaces from learned photometric priors, an ability inherited from the backbone MDE model (UniDepth) trained on large-scale datasets.
+Compared to baselines, POLAR reduces MAE by 4.4% and RMSE by 3.7% on nuScenes, 38.5% and 57.5% on ZJU, and 31.8% and 38.5% on VoD, achieving SOTA results for all datasets (Tab. 1).
+We qualitatively compare POLAR against baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (Fig.
+4).
+POLAR more accurately predicts metric scale and exhibits fewer global misalignments.
+Furthermore, it is able to extrapolate depth predictions of object surfaces from learned photometric priors, an ability inherited from the backbone MDE model (e.g., UniDepth) trained on large-scale datasets.
 Our polynomial fitting approach then non-uniformly corrects depth discrepancies across the scene, leading to significant performance improvement over initial MDE predictions.
-Introducing multiple degrees of freedom enables PolyRad to correct misalignments of objects relative to each other, improving accuracy by correcting consistent over- or under-estimations in local regions.
-For instance, the raw MDE prediction for Image C incorrectly places the bus stop roof at around the same depth as the more distant tree branch, as highlighted on the right.
+Introducing multiple degrees of freedom enables POLAR to correct misalignments of objects relative to each other, improving accuracy by correcting consistent over- or under-estimations in local regions.
+For instance, the raw MDE prediction for Image C incorrectly places the bus stop roof at around the same depth as the more distant tree branches, as highlighted on the right.
 Additionally, MDE does not accurately infer the boundary between the curb and asphalt, again placing both at erroneously similar depths.
-With a learned polynomial transformation, PolyRad corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
+With a learned polynomial transformation, POLAR corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
 While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure.
-In Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths.
+As shown in the top-left highlighted region of Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths.
 Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts.
@@
 GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
+As the polynomial degree—selected as a hyperparameter—increases, the maximum possible number of inflection points also grows, giving POLAR greater expressive power to correct misalignments between the many scene elements in Fig.
+3, including cars, trucks, trees, buildings, and traffic lights.
+This is reflected in the successively improved error maps.
+See Sec.
+A for quantitative evaluations.
+4.2 Computational Efficiency
+Table 2: Training times (minutes per epoch) on nuScenes using an NVIDIA A6000 GPU.
+Method Train Time / Epoch
+Lin (Lin et al., 2020) 89.25
+RadarNet (Singh et al., 2023) 101.50
+SparseBeatsDense (Li et al., 2024b) 63.96
+GET-UP (Sun et al., 2025) 249.57
+RadarCam-Depth (Li et al., 2024a) 86.38
+POLAR (Ours) 33.16
+Table 3: Inference time in milliseconds (ms) and computational cost (GFLOPs) on nuScenes.
+Method Inference Time GFLOPs
+Lin (Lin et al., 2020) 129.96 550.43
+SparseBeatsDense (Li et al., 2024b) 97.47 532.74
+GET-UP (Sun et al., 2025) 445.45 630.99
+RadarCam-Depth (Li et al., 2024a) 315.64 619.02
+TacoDepth (Wang et al., 2025) 29.30 139.87
+POLAR (Ours) 24.81 89.70
+POLAR achieves state-of-the-art training time, inference time (ms), and computational overhead (GFLOPs) compared to all baseline methods, while also achieving state-of-the-art accuracy.
+Tab.
+3 shows that POLAR has the lowest training time per epoch among all methods, requiring 33.16 minutes per epoch on nuScenes.
+This efficiency stems from our streamlined design, which avoids multi-stage processing and explicit radar-camera association learning, both of which contribute to longer training times in methods such as RadarCam-Depth and GET-UP.
+Tab.
+3 highlights POLAR’s state-of-the-art inference speed, requiring just 24.81 ms per frame—a reduction of 15.3% compared to the previous state-of-the-art TacoDepth and 92.1% compared to RadarCam-Depth.
+This corresponds to 40.3 fps, enabling real-time depth perception.
+Furthermore, POLAR achieves this inference speed with a lower computational overhead of 89.70 GFLOPs—a 39.5% reduction relative to TacoDepth and an 85.5% reduction relative to RadarCam-Depth.
+Taken together, these results establish POLAR as not only the most accurate radar-camera depth estimation method, but also the most practical for real-time deployment where latency is critical.
+5 Discussion
+Limitations.
+POLAR requires tuning the polynomial degree as a hyperparameter (Sec.
+A), which may vary across datasets.
+Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
+Summary.
+We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth.
+Our principled and efficient approach demonstrates that a fundamental insight—shifting from affine scale-and-shift to flexible polynomial transformations—outperforms computationally heavier methods.
+Appendix
+Appendix A Comparative Studies
+Table 4: MDE backbone comparative studies on nuScenes.
 nuScenes
 Method MAE RMSE
-DPT [41] 5188.2 6884.5
-Depth Anything [56] 2404.9 4851.1
-Depth Pro [4] 3835.0 6600.3
-UniDepth [37] 2129.8 4887.7
+DPT (Ranftl et al., 2021) 5188.2 6884.5
+Depth Anything (Yang et al., 2024) 2404.9 4851.1
+Depth Pro (Bochkovskii et al., 2024) 3835.0 6600.3
+UniDepth (Piccinelli et al., 2025) 2129.8 4887.7
 RadarCam-Depth w/ DPT 1689.7 3948.0
@@
 RadarCam-Depth w/ UniDepth 1872.0 4321.2
-PolyRad w/ DPT 1525.6 3745.0
-PolyRad w/ Depth Anything 1515.1 3719.4
-PolyRad w/ Depth Pro 1627.7 4143.9
-PolyRad w/ UniDepth 1454.1 3488.9
-Table 2: MDE backbone comparative studies on nuScenes.
-4.2 Comparative Studies
+POLAR w/ DPT 1525.6 3745.0
+POLAR w/ Depth Anything 1515.1 3719.4
+POLAR w/ Depth Pro 1627.7 4143.9
+POLAR w/ UniDepth 1407.8 3193.5
 MDE Models.
-We evaluate PolyRad using four monocular depth estimation (MDE) backbones: DPT [41], Depth Anything [56], Depth Pro [4], and UniDepth [37].
-DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see Tab.
-2 and Supp.
-Mat), and thus we process them the same way as scaleless depth.
+We evaluate POLAR using four monocular depth estimation (MDE) backbones: DPT (Ranftl et al., 2021), Depth Anything (Yang et al., 2024), Depth Pro (Bochkovskii et al., 2024), and UniDepth (Piccinelli et al., 2025).
+DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see raw MDE performance in Tabs.
+41010), and thus we process them the same way as scaleless depth.
 Tab.
-2 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and PolyRad.
+4 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and POLAR.
 Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements.
-In contrast, PolyRad consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting.
+In contrast, POLAR consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting.
 For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results.
-Among all configurations, PolyRad w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 30.2% and over RadarCam-Depth w/ UniDepth by 20.8%.
-Polynomial Degree.
-As polynomial degree—selected as a hyperparameter—increases, so does the number of inflection points, allowing PolyRad to correct misalignments between the many scene elements in Fig.
-4, including cars, trucks, trees, buildings, and traffic lights.
-This is reflected in the successively improved error maps.
-Tab.
-3 quantifies the performance gains achieved with higher-degree polynomials.
-At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
-Ablations.
-We evaluate the impact of key components of PolyRad in Tab.
-4.
-Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations.
-Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features.
-Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
+Among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 51.2% and over RadarCam-Depth w/ UniDepth by 37.8%.
+Table 5: Sensitivity study of polynomial degree, selected as a hyperparameter.
 nuScenes ZJU
@@
 6 1466.9 3496.0 670.3 1297.9
-8 1454.1 3488.9 629.6 1171.3
+8 1407.8 3193.5 629.6 1171.3
 10 1463.7 3494.5 643.3 1184.5
-Table 3: Sensitivity study of polynomial degree hyperparameter.
+Table 6: Ablation studies of architecture and loss components.
 nuScenes VoD
@@
 monotonicity loss 1921.1 4399.5 1924.5 4660.3
-no ablations 1454.1 3488.9 1500.1 3951.8
-Table 4: Ablation studies of architecture and loss components.
-5 Discussion
-Limitations.
-PolyRad requires tuning the polynomial degree as a hyperparameter, which may vary across datasets.
-Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
-Summary.
-We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth.
-Our principled and efficient approach demonstrates that a fundamental insight—shifting from simple scale-and-shift to flexible, radar-guided polynomial transformations—can outperform more complex systems.
+pos. embeddings 1454.1 3488.9 1500.1 3951.8
+no ablations 1407.8 3193.5 629.6 1171.3
+Polynomial Degree.
+Tab.
+6 quantifies the performance gains achieved with higher-degree polynomials, with degree 8 outperforming lower degrees.
+At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
+Figure 5: To quantify absolute improvement in our qualitative results, we provide the colorbars that were used in all examples in Figs. 346, and 7.
+Ablations.
+We evaluate the impact of key components of POLAR in Tab.
+6.
+Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations.
+Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features.
+Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
+Figure 6: Additional qualitative comparison on VoD.
+POLAR more accurately predicts metric scale, and exhibits fewer global misalignments and geometric artifacts in comparison to other methods.
+Polynomial fitting, allowing POLAR to non-uniformly scale distinct local regions, results in notable improvements in metric depth estimation over initial MDE predictions.
+This is apparent in all MDE error maps: though object depth geometry is reasonable, actual metric depth values tend to be inaccurate.
+GET-UP, while marginally improving metric depth estimation relative to MDE, presents substantial errors in depth structure.
+For instance, the presence of geometric artifacts or omission of objects or regions entirely can be seen through the following highlighted examples: loss of motorcycle body and visor in Image A, front portion of biker (right) and bike wheel (far-left) in Image B, top of motorcyclist in Image C, and street sign in Image D.
+RadarCam-Depth (RC-D) similarly exhibits geometric artifacts, present in the sign pole, motorcycle, and fence post in Image A, biker and bicycle wheel in Image B, roof overhang and motorcyclist in Image C, and fence posts and pedestrians in Image D.
+In contrast, POLAR extrapolates correct structure from photometric priors while also demonstrating clear improvements in metric depth fidelity compared to all other baselines.
+Appendix B More on Computational Efficiency
+Incremental increase in computation for each additional polynomial degree.
+Adding a $k+1\text{-th}$ polynomial term requires minimal computational overhead.
+For an MDE prediction of shape $(H,W)$, the additional cost consists of three components: predicting another coefficient via the linear layer of ($1\times 64\times 2$ FLOPs), computing the $(k+1)\text{-th}$ power from pre-computed $k\text{-th}$ exponentiation ($H\times W$ multiplications), and incorporating this term into the final depth prediction ($2\times H\times W$ operations for multiplication by the coefficient and addition).
+For nuScenes, where each image has shape $(900,1600)$, each additional polynomial term incurs an additional computational cost of $0.0043$ GFLOPs, which is a $0.0048\%$ increase.
+Appendix C Dataset Details
+[…]
+[…]
+[…]
+Appendix D Full nuScenes Benchmark
+Table 7: Full quantitative results ($m$) on the nuScenes benchmark.
+nuScenes
+Distance Method MAE RMSE
+50m NLSPN (Park et al., 2020) 2790.0 5813.4
+BPNet (Tang et al., 2024) 2407.0 4438.0
+RC-PDA (Long et al., 2021) 2225.0 4156.5
+RC-PDA-HG (Long et al., 2021) 2210.0 4234.0
+BTS (Lee et al., 2019) 1937.0 3885.0
+DORN (Lo & Vandewalle, 2021) 1926.6 4124.8
+RadarNet (Singh et al., 2023) 1727.7 3746.8
+CaFNet (Sun et al., 2024) 1674.0 3674.0
+Lin (Lin et al., 2020) 1598.2 3790.1
+SparseBeatsDense (Li et al., 2024b) 1524.5 3567.3
+RadarCam-Depth (Li et al., 2024a) 1286.1 2964.3
+GET-UP (Sun et al., 2025) 1241.0 2857.0
+TacoDepth (Wang et al., 2025) 1046.8 2487.5
+POLAR (Ours) 1014.4 2475.7
+70m NLSPN (Park et al., 2020) 3140.0 6580.6
+RC-PDA (Long et al., 2021) 3326.1 6700.6
+RC-PDA-HG (Long et al., 2021) 3485.6 7002.9
+BTS (Lee et al., 2019) 2346.0 4811.0
+DORN (Lo & Vandewalle, 2021) 2170.0 4532.0
+RadarNet (Singh et al., 2023) 2073.2 4590.7
+CaFNet (Sun et al., 2024) 2010.0 4493.0
+Lin (Lin et al., 2020) 1897.8 4558.7
+SparseBeatsDense (Li et al., 2024b) 1822.9 4303.6
+RadarCam-Depth (Li et al., 2024a) 1587.9 3662.5
+GET-UP (Sun et al., 2025) 1541.0 3657.0
+TacoDepth (Wang et al., 2025) 1347.1 3152.8
+POLAR (Ours) 1286.1 2947.3
+80m NLSPN (Park et al., 2020) 3257.7 6872.4
+RC-PDA (Long et al., 2021) 3721.0 7632.0
+RC-PDA-HG (Long et al., 2021) 3664.0 7775.0
+AdaBins (Bhat et al., 2021) 3541.0 5885.0
+P3Depth (Patil et al., 2022) 3130.0 5838.0
+LapDepth (Song et al., 2021) 2544.0 5151.0
+PnP (Wang et al., 2018) 2496.0 5578.0
+BTS (Lee et al., 2019) 2467.0 5125.0
+DORN (Lo & Vandewalle, 2021) 2432.0 5304.0
+RadarNet (Singh et al., 2023) 2179.3 4898.7
+CaFNet (Sun et al., 2024) 2109.0 4765.0
+Lin (Lin et al., 2020) 1988.4 4841.1
+SparseBeatsDense (Li et al., 2024b) 1927.0 4609.6
+RadarCam-Depth (Li et al., 2024a) 1689.7 3948.0
+GET-UP (Sun et al., 2025) 1632.0 3932.0
+TacoDepth (Wang et al., 2025) 1492.4 3324.8
+POLAR (Ours) 1407.8 3193.5
+We present the full set of quantitative results on the nuScenes dataset in Tab.
+7.
+This table includes additional baseline methods that were omitted from the main text for brevity, providing a comprehensive comparison of POLAR against all known existing radar-camera depth estimation methods.
+As shown, POLAR consistently outperforms all competing methods across all maximum distance thresholds (50m, 70m, and 80m), achieving the lowest mean absolute error (MAE) and root mean squared error (RMSE).
+Appendix E Comparison to Depth Completion
+One potential idea for radar-camera depth estimation is to apply lidar-camera depth completion methods designed to densify sparse depth maps using surrounding context.
+One such method, BPNet (Tang et al., 2024) achieves state-of-the-art performance on the KITTI depth completion benchmark by leveraging bilateral propagation.
+However, when applied to radar-camera depth estimation, BPNet performs poorly, as shown in Tab.
+7.
+POLAR outperforms BPNet by 57.9% in MAE and 44.2% in RMSE on nuScenes, highlighting the limitations of directly applying lidar depth completion methods to radar data.
+The key reason for this underperformance lies in the fundamental differences between lidar and radar point clouds.
+Unlike lidar, radar measurements are orders of magnitude sparser and significantly noisier due to factors such as limited antenna elements, elevation ambiguity (see Fig.
+7), and lower range resolution (Singh et al., 2023).
+Lidar depth completion methods assume a relatively dense and structured input (Xia et al., 2023), leveraging local spatial continuity to propagate depth estimates effectively.
+In contrast, radar points are too sparse for such methods to infer meaningful local depth relationships, leading to poor depth reconstruction when attempting direct densification.
+Table 8: Error metrics for depth estimation. These metrics compute the error between predicted depth $\hat{d}(x)$ and ground truth depth $d(x)$.
+Metric Definition
+MAE $\downarrow$ $\frac{1}{|\Omega|}\sum_{x\in\Omega}|\hat{d}(x)-d(x)|$
+RMSE $\downarrow$ $\Big(\tfrac{1}{|\Omega|}\sum_{x\in\Omega}|\hat{d}(x)-d(x)|^{2}\Big)^{1/2}$
+In addition, we compare against Non-Local Spatial Propagation Network (NLSPN) (Park et al., 2020), which achieves near state-of-the-art performance on the KITTI depth completion benchmark by refining sparse lidar depth with an iterative non-local spatial propagation procedure.
+NLSPN predicts an initial depth map along with pixel-wise confidences, then refines it by estimating non-local neighbors and their corresponding affinities to selectively propagate depth information.
+Unlike other approaches that rely on fixed local neighbors, NLSPN adaptively determines relevant non-local neighbors, improving depth completion accuracy, especially near depth boundaries.
+However, again, when applied to radar-camera depth estimation, NLSPN performs poorly, as shown in Tab.
+7.
+POLAR outperforms NLSPN by 59.8% in MAE and 55.4% in RMSE on nuScenes.
+Instead of relying on depth completion-style densification, POLAR directly learns a transformation function from radar to metric depth by leveraging polynomial fitting.
+Our method avoids the pitfalls of propagating unreliable local depth information by refining MDE predictions with learned polynomial coefficients, enabling flexible, scene-adaptive depth corrections that effectively capture object relationships and global scene structure.
+Figure 7: nuScenes dataset visualization.
+The elevation ambiguity of radar points results in erroneous projection onto the image plane that makes it challenging for depth completion methods to infer dense depth.
+In contrast, lidar points yield a denser, image-aligned projection, which is why accurate 3D scene reconstruction with depth completion methods is possible.
+Appendix F Evaluation Metrics and Implementation Details
+The evaluation metrics used in our study, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), are formulated in Tab.
+8.
+Lower values equal better performance for both MAE and RMSE.
+Unless specified otherwise, all reported values are in millimeters (mm).
+We train for 60 epochs using a cosine decay learning rate scheduler with learning rate of $5\times 10^{-5}$, and use weighting terms $\lambda_{L_{1}}=1.0,\lambda_{L_{2}}=0.4,\lambda_{mono}=0.25$ for our loss function.
+Appendix G Additional Comparisons
+Full MDE Comparisons.
+For ZJU (see Tab.
+10), among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 61.1% and over RadarCam-Depth w/ UniDepth by 54.2%.
+For VoD (see Tab.
+10), POLAR w/ UniDepth achieves the best performance in MAE, improving over raw UniDepth predictions by 42.4% and over RadarCam-Depth w/ UniDepth by 32.7%, while POLAR w/ Depth Anything achieves the best performance in RMSE, improving over RadarCam-Depth w/ Depth Anything by 37.6% and over raw Depth Anything predictions (inverted and median scaled) by 10.4%.
+Table 9: MDE backbone comparative studies on ZJU-4DRadarCam (ZJU).
+ZJU
+Method MAE RMSE
+DPT (Ranftl et al., 2021) 1885.3 3326.1
+Depth Anything (Yang et al., 2024) 1943.2 3469.3
+Depth Pro (Bochkovskii et al., 2024) 1680.2 3144.9
+UniDepth (Piccinelli et al., 2025) 1533.0 3188.4
+RadarCam-Depth w/ DPT 1183.5 3229.0
+RadarCam-Depth w/ Depth Anything 1724.4 3661.3
+RadarCam-Depth w/ Depth Pro 1490.6 3429.5
+RadarCam-Depth w/ UniDepth 1152.5 3168.6
+POLAR w/ DPT 707.1 1216.9
+POLAR w/ Depth Anything 657.2 1225.4
+POLAR w/ Depth Pro 640.3 1174.8
+POLAR w/ UniDepth 629.6 1171.3
+Table 10: MDE backbone comparative studies on View-of-Delft.
+View-of-Delft
+Method MAE RMSE
+DPT (Ranftl et al., 2021) 4117.9 5498.9
+Depth Anything (Yang et al., 2024) 3270.5 4411.9
+Depth Pro (Bochkovskii et al., 2024) 3275.9 5936.7
+UniDepth (Piccinelli et al., 2025) 2605.6 5691.0
+RadarCam-Depth w/ DPT 4013.5 5911.9
+RadarCam-Depth w/ Depth Anything 3103.6 6328.7
+RadarCam-Depth w/ Depth Pro 2843.4 6082.0
+RadarCam-Depth w/ UniDepth 2227.4 5385.8
+POLAR w/ DPT 1891.4 4252.6
+POLAR w/ Depth Anything 1770.5 3951.8
+POLAR w/ Depth Pro 1520.2 3987.2
+POLAR w/ UniDepth 1500.1 3960.5
+Leveraging Radar.
+Tab.
+13 shows that replacing radar points with learnable, dataset-specific points worsens MAE by 28.0% and RMSE by 29.9%, demonstrating that we indeed leverage the radar inputs effectively.
+As additional evidence, Tab.
+12 shows that our method, evaluated zero-shot cross-dataset, achieves comparable or better performance than baselines trained on the target datasets.
+Tab.
+12 shows we are more robust to reduced radar point density at inference, i.e., less performance degradation than the baseline method RadarCam-Depth.
+Table 11: Zero-shot generalization.
+$\rightarrowZJU nuScenes\rightarrowVoD$
+Method MAE RMSE MAE RMSE
+GET-UP (zero-shot) 3845.2 8469.7 4809.1 8653.9
+RadarCam-Depth (zero-shot) 5435.9 9785.8 7521.5 9194.8
+GET-UP (trained) 1699.7 3882.6 2917.3 6145.1
+RadarCam-Depth (trained) 1183.5 3229.0 2227.4 5385.8
+Ours (zero-shot) 1147.9 3109.5 2256.2 4744.2
+Table 12: Reduced radar point density.
+RadarCam-Depth Ours
+% radar kept / removed MAE RMSE MAE RMSE
+25% kept / 75% removed 7969.4 10831.5 2416.4 4836.6
+50% kept / 50% removed 4819.8 7077.7 1816.8 3945.7
+75% kept / 25% removed 2537.3 4247.4 1575.2 3611.8
+100% kept / 0% removed 1689.7 3948.0 1407.8 3193.5
+Table 13: Additional comparisons of POLAR vs. learnable dataset-specific points in place of radar points, and regression baselines.
+MAE RMSE
+Learnable Points 1860.9 4207.1
+Isotonic Reg. 2895.2 4340.5
+Cubic Hermite 2131.3 4588.0
+PCHIP 1809.6 4054.7
+LoRA 2030.6 4493.7
+ViT-Adapter 1859.6 4280.0
+Our Performance 1407.8 3193.5
+Adapters.
+Tab.
+13 shows that recent adapter-based finetuning methods LoRA (Hu et al., 2022) and ViT-Adapter (Chen et al., 2022), even with a post-hoc linear fit to projected radar points, do not outperform us.
+Regression Baselines.
+Isotonic regression and monotone spline fitting methods are natural baselines.
+Tab.
+13 shows these methods for regressing projected radar points on MDE predictions do not outperform us.
+We hypothesize that this is due to noise in radar points that can be mitigated by learning (Sec.
+3.2).
+Our learned polynomial fit may, in principle, introduce unwanted inversions of initially correct MDE predictions.
+POLAR successfully mitigates this effect through the proposed novel first-derivative regularization term (see Sec.
+3.4 and Eqs.
+67), which effectively constrains such inversions.
+To quantify this, we compute Kendall’s $\tau$ coefficient between predicted and ground-truth depths.
+Our method achieves the highest monotonicity ($\tau=0.969$) over regression baselines Isotonic Regression ($\tau=0.871$), PCHIP ($\tau=0.758$), and Cubic Hermite Spline ($\tau=0.736$).
+The raw MDE predictions do exhibit monotonicity with respect to ground truth ($\tau=0.957$), but our polynomial transformation increases it, indicating that we correct unwanted inversions.
+To assess statistical significance, we compute Kendall’s $\tau$ over 30 bootstrap samples for both our method and the raw MDE predictions.
+A two-sample t-test reveals a statistically significant difference in mean monotonicity ($p=0.012$).
+Appendix H Proof: Limitations of Global Scale and Shift
+As further theoretical justification, we prove by construction that an affine scale-and-shift transformation is insufficient to fit MDE predictions to ground truth.
+Proposition 1.
+There exist infinitely many sets of $k\geq 3$ (MDE prediction $\hat{d}$, ground truth $d$) pairs such that no global scale $\alpha$ and shift $\beta$ satisfy $d=\alpha\hat{d}+\beta$ for all $k$ pairs simultaneously.
+Proof by Construction.
+Consider the following three (MDE prediction $\hat{d}$, ground truth $d$) pairs from the nuScenes dataset, specifically from the image shown in Fig. 3:
+$(\hat{d},d)\in\{(5,7),(10,9),(15,44)\}.$
+Assume to the contrary that there exist $\alpha,\beta\in\mathbb{R}$ such that
+$d=\alpha\hat{d}+\beta$
+holds for all pairs.
+From the first two pairs, we obtain:
+$7=5\alpha+\beta,\quad 9=10\alpha+\beta.$
+Subtracting gives $\alpha=0.4$ and $\beta=5$ as the unique solution for these two pairs.
+Applying this solution to the third pair yields:
+$\alpha\cdot 15+\beta=0.4\cdot 15+5=11,$
+which contradicts the required equality with the ground truth value $d=44$, since the residual error equals 4411=3344-11=33 and not zero.
+Hence no global scale $\alpha$ and shift $\beta$ exist that can satisfy all three pairs simultaneously.
+Moreover, scaling each pair by any nonzero constant produces infinitely many distinct $3$-sets of ($\hat{d},d$) pairs for which no $\alpha$ and $\beta$ exists.
+Then, for any such $3$-set, appending k3k-3 arbitrary pairs yields an infinite family of $k$-sets ($k>3$) that likewise admit no solution.
+Corollary 1.
+It is therefore a misconception that the scale ambiguity in MDE can be resolved solely by a global scale and shift.
+In contrast, any smooth relationship between $\hat{d}$ and $d$ can be locally approximated by a polynomial via Taylor expansion, giving our polynomial fitting formulation the theoretical capacity to approximate $d$ as a function of $\hat{d}$ arbitrarily well.

### v3

@@
 Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a fundamental insight: although MDE models often infer reasonable local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale and shift (affine) transformation insufficient given three or more of these regions.
-To address this limitation, we use polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges.
+To address this limitation, we use polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust predictions non-uniformly across depth ranges.
 In this way, POLAR generalizes beyond affine transformations and is able to correct such misalignments by introducing inflection points.
@@
 1 Introduction
-Metric 3D reconstruction is critical for spatial tasks such as self-driving (Maier et al., 2012; Gupta et al., 2021), where it is necessary for one to perceive the structure of the 3D environment in order to navigate.
+Metric 3D reconstruction is critical for spatial tasks such as self-driving [maier2012real, GUPTA2021100057], where it is necessary for one to perceive the structure of the 3D environment in order to navigate.
 In many such systems, multiple sensors—including cameras, lidar, and radar—provide complementary information about the 3D scene.
-While lidar sensors offer dense and precise point clouds, they are expensive and not widely available (Raj et al., 2020).
-In contrast, radar sensors, particularly millimeter wave (mmWave) radars (Iizuka et al., 2003), return only about a hundred points per frame and are noisier (Han et al., 2023).
-Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles (Eichelberger & McCartt, 2016).
+While lidar sensors offer dense and precise point clouds, they are expensive and not widely available [raj2020survey].
+In contrast, radar sensors, particularly millimeter wave (mmWave) radars [iizuka2003millimeter], return only about a hundred points per frame and are noisier [han20234d].
+Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles [eichelberger2016toyota].
+Figure 1: If an MDE model predicts incorrect relative depths between three or more objects, an affine scale-and-shift (dashed) cannot resolve this misalignment.
+POLAR (solid) overcomes this limitation by learning an NN’th-order polynomial fit with up to N2N-2 inflection points.
 Unless camera baselines are known, images offer scaleless reconstruction.
@@
 Training robust models that can generalize across these factors often demands large-scale datasets, which are costly to collect.
-As a result, leveraging pretrained monocular depth estimation (MDE) foundation models (Ranftl et al., 2021; Yang et al., 2024; Bochkovskii et al., 2024; Piccinelli et al., 2025) emerges as a practical alternative.
-However, as monocular 3D reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or depth that lacks the fidelity needed for applications demanding accurate metric-scale reconstruction, such as mapping and navigation.
-Existing approaches (Yin et al., 2023; Viola et al., 2024; Hu et al., 2024; Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025) attempt to transform these MDE predictions into accurate metric depth maps using a global affine (scale-and-shift) transformation.
+As a result, leveraging pretrained monocular depth estimation (MDE) foundation models [ranftl2021vision, depth_anything_v2, depth-pro, piccinelli2025unidepthv2] emerges as a practical alternative.
+However, as 3D reconstruction from a single 2D image is inherently ill-posed, they typically infer scaleless relative depth, or depth that lacks the fidelity needed for applications demanding accurate metric-scale reconstruction, such as mapping and navigation.
+Existing approaches [yin2023metric3d, viola2024marigold, hu2024depthcrafter, zeng2025rsa, ding2025fixing, yu2025relative] attempt to transform these MDE predictions into accurate metric depth maps using a global affine (scale-and-shift) transformation.
 This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene.
@@
 1).
-Figure 1: If an MDE model predicts incorrect relative depths between three or more objects, an affine scale-and-shift (dashed) cannot resolve this misalignment.
-POLAR (solid) overcomes this limitation by learning an NN’th-order polynomial fit with up to N2N-2 inflection points.
 In this work, we challenge the common assumption that scale ambiguity in MDE is only up to an unknown global scale and shift.
@@
 This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and cross-region misalignments.
-While one may adjust every prediction (pixel) of an MDE model (Li et al., 2024a), the number of degrees of freedom is the total number of pixels.
+While one may adjust every prediction (pixel) of an MDE model [li2024radarcam], the number of degrees of freedom is the total number of pixels.
 Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem.
@@
 Monocular Depth Estimation.
-With recent advances in the scalability of neural networks (Dosovitskiy et al., 2021; Caron et al., 2021), monocular depth estimation (MDE) models can infer scaleless relative depth from a single image across diverse, unseen domains (Ranftl et al., 2020; Yang et al., 2024; Piccinelli et al., 2025; Bochkovskii et al., 2024), benefiting from large-scale datasets.
-Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation (Hoyer et al., 2023; Bian et al., 2019), uncertainty estimation (Poggi et al., 2020), optical flow (Zhao et al., 2020), and visual odometry (Song et al., 2023), to improve relative depth estimation.
-However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem.
-Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity (Viola et al., 2024).
-One way to address this limitation is to incorporate range-sensing modalities such as lidar (Jaritz et al., 2018; Ezhov et al., 2024), radar (Singh et al., 2023), or visual-inertial odometry (Wong et al., 2020).
+With recent advances in the scalability of neural networks [dosovitskiy2020image, caron2021emerging], monocular depth estimation (MDE) models can infer scaleless relative depth from a single image across diverse, unseen domains [ranftl2020towards, depth_anything_v2, piccinelli2025unidepthv2, depth-pro], benefiting from large-scale datasets.
+Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation [hoyer2023improving, bian2019unsupervised], uncertainty estimation [poggi2020uncertainty], optical flow [zhao2020towards], and visual odometry [song2023unsupervised], to improve relative depth estimation.
+However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating depth from a single image is an ill-posed problem.
+Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity [viola2024marigold].
+One way to address this limitation is to incorporate range-sensing modalities such as lidar [jaritz2018sparse, ezhov2024all], radar [singh2023depth], or visual-inertial odometry [wong2020unsupervised].
 Image-Guided Depth Completion.
 Most often studied in the context of lidar-camera depth estimation, image-guided depth completion leverages the strengths of each modality: images offer dense visual context and structural priors, while lidar points provide absolute metric scale to resolve depth ambiguity.
-Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers (Ma & Karaman, 2018; Ma et al., 2018), late fusion, where inputs are processed by independent branches (Yan et al., 2021; Rim et al., 2025), and multi-scale fusion, which captures both local details and global scene structure (Li et al., 2020).
-U-Net-like architectures have been widely used for coarse-to-fine depth completion (Hu et al., 2021; Lin et al., 2022), with improvements from deformable convolutions (Park et al., 2020; Xu et al., 2020), and attention mechanisms (Rho et al., 2022; Zhang et al., 2023).
+Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers [sparse_to_dense, Ma2018SelfSupervisedSS], late fusion, where inputs are processed by independent branches [RigNet, rim2025protodepth], and multi-scale fusion, which captures both local details and global scene structure [msg_chn].
+U-Net-like architectures have been widely used for coarse-to-fine depth completion [Hu2021PENetTP, Lin2022DynamicSP, chen2024uncle], with improvements from deformable convolutions [Park2020NonLocalSP, Xu2020DeformableSP], and attention mechanisms [Rho2022GuideFormerTF, CompletionFormer].
 Radar-Camera Depth Estimation.
-While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions (Raj et al., 2020).
-In contrast, radar provides a cost-effective alternative (Hunt et al., 2024), offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle (Paek et al., 2022; Srivastav & Mandal, 2023).
-The ubiquity of radar sensors (Eichelberger & McCartt, 2016) in existing automotive and robotic platforms further supports their integration into depth estimation pipelines.
+While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions [raj2020survey].
+In contrast, radar provides a cost-effective alternative [hunt2024radcloud], offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle [paek2022k, srivastav2023radars].
+The ubiquity of radar sensors [eichelberger2016toyota] in existing automotive and robotic platforms further supports their integration into depth estimation pipelines.
 Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
-Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity (Singh et al., 2023) of radar point clouds.
-(Lin et al., 2020) uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map.
-(Long et al., 2021) leverages Doppler velocity and optical flow to associate radar points with image pixels.
-(Lo & Vandewalle, 2021) refines radar depth using height-extension of radar points to address elevation ambiguity.
-(Singh et al., 2023) introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion.
-(Li et al., 2024b) mitigates distribution artifacts using sparse supervision, while (Sun et al., 2024) employs a two-stage confidence-driven approach.
-GET-UP (Sun et al., 2025) uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features.
-Like our method, RadarCam-Depth (Li et al., 2024a) and TacoDepth (Wang et al., 2025) also begin with an initial MDE prediction, but refine it and decode dense depth directly from fused features, rather than employing a learned scene-fitting approach.
+Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity [singh2023depth] of radar point clouds.
+[lin2020depth] uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map.
+[long2021full] leverages Doppler velocity and optical flow to associate radar points with image pixels.
+[lo2021depth] refines radar depth using height-extension of radar points to address elevation ambiguity.
+[singh2023depth] introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion.
+[li2023sparse] mitigates distribution artifacts using sparse supervision, while [sun2024cafnet] employs a two-stage confidence-driven approach.
+GET-UP [sun2024get] uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features.
+Like our method, RadarCam-Depth [li2024radarcam] and TacoDepth [wang2025tacodepth] also begin with an initial MDE prediction, but refine it and decode dense depth directly from fused features, rather than employing a learned scene-fitting approach.
 Existing methods operate within the paradigm of completion or direct decoding of fused features, and they often rely on multi-stage training and explicit radar-pixel association learning, increasing model complexity and computational overhead.
 In contrast, POLAR employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale.
-It bypasses the need for multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
+It bypasses the need for multi-stage systems that learn explicit correspondences while enabling flexible depth corrections.
+Figure 2: Method Overview.
+POLAR transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features.
+Learnable prototypes extract patterns in the configurations of radar point clouds and are used to aggregate spatially-informed radar features.
+The geometry-aware MDE features are fused with the radar features via a learnable soft-correspondence module to yield a unified scene representation that is used to predict polynomial coefficients for fitting.
+This enables non-uniform corrections that improve accuracy beyond affine transformations.
 3 Method Formulation
@@
 Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$.
-Figure 2: Method Overview.
-POLAR transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features.
-Learnable prototypes extract patterns in the configurations of radar point clouds and are used to aggregate spatially-informed radar features.
-The geometry-aware MDE features are fused with the radar features via a learnable soft-correspondence module to yield a unified scene representation that is used to predict polynomial coefficients for fitting.
-This enables non-uniform corrections that improve accuracy beyond affine transformations.
 3.1 Motivation
-Existing methods (Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025; Viola et al., 2024; Hu et al., 2024; Yin et al., 2023) refine MDE predictions by applying simple scale-and-shift transformations.
+Existing methods [zeng2025rsa, ding2025fixing, yu2025relative, viola2024marigold, hu2024depthcrafter, yin2023metric3d] refine MDE predictions by applying simple scale-and-shift transformations.
 However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths.
@@
 1).
-This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
+This violates the assumption that MDE reconstruction is up to an unknown global scale and shift.
 To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map via summation.
@@
 Radar point clouds, while noisy and sparse, provide metric depth measurements.
-We concatenate a sinusoidal 3D positional embedding $\phi_{3D}(x,y,z)$ to $C$, and feed the resulting representation into a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
+We concatenate a sinusoidal 3D positional embedding $\phi_{3D}(x,y,z)$ to $C$, and feed it into a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
 Radar Aggregation.
@@
 Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes.
-Unlike existing works (Singh et al., 2023; Li et al., 2024a) that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud.
+Unlike existing works [singh2023depth, li2024radarcam] that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud.
 This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
-Next, learnable prototypes are matched to the most relevant radar features for a given input.
+Next, learnable prototypes are matched to the most relevant radar features.
 We treat the prototypes as centroids and perform soft clustering over the radar features.
 Each radar point is softly assigned to prototypes according to feature similarity, and the corresponding values are aggregated to yield a global scene-descriptive representation $F_{R}$:
-$D_{ij}=\|P_{j}-\Phi_{r}(F_{r})_{i}\|^{2},\quad F_{R}=\text{softmax}\!\left(-D\,/\,\tau\right)\,\Psi_{r}(F_{r}),$ (2)
-where $\Phi_{r}$, $\Psi_{r}$ are MLPs that project the radar features $F_{r}$, the matrix $D\in\mathbb{R}^{N_{P}\times N_{C}}$ contains the pairwise squared distances between prototypes $\{P_{j}\}_{j=1}^{N_{P}}$ and projected radar features $\{\Phi_{r}(F_{r})_{i}\}_{i=1}^{N_{C}}$, and $\tau$ denotes temperature.
+$D_{ij}=\|P_{j}-\Phi_{r}(F_{r})_{i}\|^{2},\quad F_{R}=\sigma\!\left(-D\,/\,\tau\right)\,\Psi_{r}(F_{r}),$ (2)
+where $\Phi_{r}$, $\Psi_{r}$ are MLPs that project the radar features $F_{r}$, the matrix $D\in\mathbb{R}^{N_{P}\times N_{C}}$ contains the pairwise squared distances between prototypes $\{P_{j}\}_{j=1}^{N_{P}}$ and projected radar features $\{\Phi_{r}(F_{r})_{i}\}_{i=1}^{N_{C}}$, and $\sigma$ denotes softmax.
 This clustering-based formulation enables prototypes to capture recurring spatial and geometric patterns in radar point clouds.
-Comparatively, applying lidar depth completion methods, which densify a sparse projection of points using surrounding context, to a radar point cloud results in poor performance (see Sec.
-E).
-This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution (Singh et al., 2023).
-MDE-Radar Fusion.
-The scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded with a learnable depth encoder $f_{z}$, producing depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$.
-These features inherit invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—and therefore primarily encode object-level properties.
-Since variations in the shape of objects and their geometry are generally more stable across scenes than photometric appearance, the resulting depth features provide a more reliable geometric context for fusion with radar features.
-This allows radar point configurations to be matched to consistently observed shapes and surfaces, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
-To this end, we learn soft spatial correspondences between the depth features $Z$ and the radar features $F_{R}$ to construct a unified scene representation $S$ that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features:
-$S=\text{softmax}\left(\frac{(Z+E)\times(\Phi_{R}(F_{R}))^{T}}{\sqrt{c_{r}}}\right)\Psi_{R}(F_{R}),$ (3)
-where $E\in\mathbb{R}^{(H\times W)\times c_{z}}$ is a learned 2D positional embedding, and $\Phi_{R}$, $\Psi_{R}$ are MLPs that project the aggregated radar features $F_{R}$.
-$c_{z}$ is set equal to $c_{r}$ to align the depth and radar features within a common embedding space, facilitating cross-modal fusion, and $S$ is reshaped to be in $\mathbb{R}^{H\times W\times c_{z}}$.
-Predicting Coefficients.
-The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield $\bar{S}=\text{GAP}(f_{s}(S))$, which is a $c_{s}$-dimensional feature vector.
-The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context.
-This final scene representation $\bar{S}$ is then fed into an MLP $\psi_{s}$, which predicts the $N+1$ polynomial coefficients as a vector $\hat{c}$:
-$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (4)
-These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities.
-Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and cross-region corrections, resulting in a high-fidelity metric-scale depth map.
-3.3 Polynomial Fitting
-Given a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ predicted by the MDE foundation model, our goal is to transform it into a metric depth map $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ using the polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\dots,\hat{c}_{N}\}$ predicted by our network (Sec.
-3.2).
-Formally, we express the final depth map $\hat{d}$ as:
-$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W.\end{split}$ (5)
-Figure 3: POLAR leverages spatial information from radar points to predict higher-degree polynomial transformations that can correct non-affine errors in MDE predictions.
-This polynomial formulation applies the learned coefficients to successive powers of the scaleless depth map, enabling complex, non-linear transformations that surpass the limitations of affine transformations.
-Each pixel’s final depth is thus computed by summing multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
-This is in contrast to directly mapping to a metric-scale depth map (Singh et al., 2023; Li et al., 2024a), where the degrees of freedom comprise every predicted pixel.
-Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system.
-As the radar point cloud ($\sim 10^{2}$ points) only occupies a small subset of the image space ($\sim 10^{6}$ pixels), the solution is underdetermined and often irregular (see Fig.
-4).
-Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab.
-6).
-POLAR offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
-In terms of overhead, compared to predicting two coefficients (scale and shift) for an affine transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost ($0.01\%$ more FLOPs): the final MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Sec.
-B for the complete derivation).
-Geometric Intuition.
-Applying polynomial fitting to the scaleless depth map is best viewed as a flexible, depth-dependent correction mechanism.
-In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
-By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates $z$.
-Lower-order coefficients ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) establish a broad global scale, providing a metric basis for the initially scaleless $z$.
-Higher-order coefficients ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) introduce both low-frequency corrections, e.g., resolving cross-object misalignments (Fig.
-1), as well as high-frequency corrections, e.g., sharpening fine object boundaries (Fig.
-4, Image C).
-The concept of inflection points provides an intuitive understanding of this approach.
-Each additional polynomial order introduces more potential inflection points, allowing the depth transformation to shift curvature where needed.
-This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction.
-Higher-order terms magnify small (i.e., high-frequency) discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct.
-In this way, the degree of the polynomial, selected as a hyperparameter, governs the capacity of our model to apply non-uniform corrections across the scene.
-The sign of the predicted polynomial coefficients furthers our interpretation.
-Positive coefficients for higher-order terms push depth values outward while negative ones pull them inward, effectively expanding or contracting selected depth intervals to correct local errors.
-In doing so, these coefficients shape the curvature of the polynomial, dictating where and how inflection points arise.
-For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations.
-By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (e.g., Fig.
-3) for refining scaleless depth into high-fidelity metric depth.
-Table 1: Quantitative results on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft test sets evaluated with various maximum evaluation distances.
+Comparatively, applying lidar depth completion methods, which densify a sparse projection of points using surrounding context, to a radar point cloud results in poor performance (see Supp.
+Mat.).
+This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution [singh2023depth].
 nuScenes ZJU View-of-Delft
@@
 POLAR (Ours) 1407.8 3193.5 629.6 1171.3 1500.1 3951.8
+Table 1: Quantitative results on nuScenes, ZJU-4DRadarCam, and View-of-Delft test sets evaluated with various maximum depths.
+MDE-Radar Fusion.
+The scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded with a learnable depth encoder $f_{z}$, producing depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$.
+These features inherit invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—and therefore primarily encode object-level properties.
+Since variations in the shape of objects and their geometry are generally more stable across scenes than photometric appearance, the resulting depth features provide a more reliable geometric context for fusion with radar features.
+This allows radar point configurations to be matched to consistently observed shapes and surfaces, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
+To this end, we learn soft spatial correspondences between the depth features $Z$ and the radar features $F_{R}$ to construct a unified scene representation $S$ that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features:
+$S=\text{softmax}\left(\frac{(Z+E)\times(\Phi_{R}(F_{R}))^{T}}{\sqrt{c_{r}}}\right)\Psi_{R}(F_{R}),$ (3)
+where $E\in\mathbb{R}^{(H\times W)\times c_{z}}$ is a learned 2D positional embedding, and $\Phi_{R}$, $\Psi_{R}$ are MLPs that project the aggregated radar features $F_{R}$.
+$c_{z}$ is set equal to $c_{r}$ to align the depth and radar features within a common embedding space, facilitating cross-modal fusion, and $S$ is reshaped to $H\times W\times c_{z}$.
+Predicting Coefficients.
+The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield $\bar{S}=\text{GAP}(f_{s}(S))$, which is a $c_{s}$-dimensional feature vector.
+The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context.
+This final scene representation $\bar{S}$ is then fed into an MLP $\psi_{s}$, which predicts the $N+1$ polynomial coefficients as a vector $\hat{c}$:
+$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (4)
+These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities.
+Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and cross-region corrections, resulting in a high-fidelity metric-scale depth map.
+Figure 3: Qualitative results on nuScenes.
+GET-UP and RadarCam-Depth (RC-D) fail to reconstruct entire regions, yielding objects with large depth errors.
+Raw MDE yields reasonable relative reconstructions but suffers from incorrect global scale and cross-object misalignments.
+POLAR leverages polynomial fitting to recover a global scale and correct these misalignments.
+3.3 Polynomial Fitting
+Given a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ predicted by the MDE foundation model, our goal is to transform it into a metric depth map $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ using the polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\dots,\hat{c}_{N}\}$ predicted by our network (Sec.
+3.2).
+Formally, we express the final depth map $\hat{d}$ as:
+$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W.\end{split}$ (5)
+This polynomial formulation applies the learned coefficients to successive powers of the scaleless depth map, enabling complex, non-linear transformations that surpass the limitations of affine transformations.
+Each pixel’s final depth is thus computed by summing multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
+This contrasts with directly mapping to a metric-scale depth map [singh2023depth, li2024radarcam], where the degrees of freedom comprise every predicted pixel.
+Each prediction of corresponding points and pixels can be viewed as a constraint or equation in our estimation system.
+As the radar point cloud ($\sim 10^{2}$ points) only occupies a small subset of the image space ($\sim 10^{6}$ pixels), the solution is underdetermined and often irregular (see Fig.
+3).
+Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab.
+4).
+POLAR offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
+In terms of overhead, compared to predicting the scale and shift for an affine transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost ($0.01\%$ more FLOPs): the final MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Supp.
+Mat.
+for the complete derivation).
+Geometric Intuition.
+Applying polynomial fitting to the scaleless depth map is best viewed as a flexible, depth-dependent correction mechanism.
+In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
+By introducing polynomial terms, our method adjusts depth according to the initial depth estimates $z$.
+Lower-order coefficients ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) establish a broad global scale, providing a metric basis for the initially scaleless $z$.
+Higher-order coefficients ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) introduce both low-frequency corrections, e.g., resolving cross-object misalignments (Fig.
+1), as well as high-frequency corrections, e.g., sharpening fine object boundaries (Fig.
+3, Image C).
+The concept of inflection points provides an intuitive understanding of this approach.
+Each additional polynomial order introduces more potential inflection points, allowing the depth transformation to shift curvature where needed.
+This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction.
+Higher-order terms magnify small (i.e., high-frequency) discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct.
+In this way, the degree of the polynomial, selected as a hyperparameter, governs the capacity of our model to apply non-uniform corrections across the scene.
+The sign of the predicted polynomial coefficients furthers our interpretation.
+Positive coefficients for higher-order terms push depth values outward while negative ones pull them inward, effectively expanding or contracting selected depth intervals to correct local errors.
+In doing so, these coefficients shape the curvature of the polynomial, dictating where and how inflection points arise.
+For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations.
+By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (e.g., Fig.
+4) for refining scaleless depth into high-fidelity metric depth.
+Figure 4: POLAR leverages spatial information from radar points to predict higher-degree polynomial transformations that can correct non-affine errors in MDE predictions.
 3.4 Loss Function
 We employ a loss function comprising three terms, each weighted by its $\lambda$, to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
-$\mathcal{L}=\lambda_{L_{1}}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{L_{2}}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{\text{mono}}\left\|\mathbf{1}_{H\times W}-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (6)
-where $d_{\text{gt}}\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
+$\mathcal{L}=\lambda_{1}\|\hat{d}-d\|_{1}+\lambda_{2}\|\hat{d}-d\|_{2}^{2}+\lambda_{\text{m}}\left\|\mathbf{1}_{H\times W}-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (6)
+where $d\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
 The first two terms, $L_{1}$ and $L_{2}$ losses, ensure that our predicted depth map $\hat{d}$ closely matches the ground truth $d_{\text{gt}}$ by penalizing discrepancies.
@@
 $\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}(x,y)=\sum_{i=1}^{N}\,i\,\hat{c}_{i}\,z(x,y)^{i-1}.$ (7)
-This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression (Barlow et al., 1972).
+This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression [barlow1972isotonic].
 Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth.
@@
 Our regularization term addresses this by preserving local depth ordinality while still allowing corrections of cross-object misalignments.
-This constraint prevents overfitting of nonlinear transformations to noisy radar data, spurious correlations, and outliers, thereby avoiding the oscillatory behavior that is a known challenge with fitting higher-degree polynomials (Bishop, 2006).
-Figure 4: Qualitative results on nuScenes.
-GET-UP and RadarCam-Depth (RC-D) fail to reconstruct entire regions, yielding objects with large depth errors.
-Raw MDE yields reasonable relative reconstructions but suffers from incorrect global scale and cross-object misalignments.
-POLAR leverages polynomial fitting to recover a global scale and correct these misalignments.
-See Fig.
-5 for colorbars.
+This constraint prevents overfitting of nonlinear transformations to noisy radar data, spurious correlations, and outliers, thereby avoiding the oscillatory behavior that is a known challenge with fitting higher-degree polynomials [bishop2006pattern].
 4 Experiments
 Baseline Methods.
-We consider five recent radar-camera depth estimation baselines: RadarNet (Singh et al., 2023), SparseBeatsDense (Li et al., 2024b), GET-UP (Sun et al., 2025), RadarCam-Depth (Li et al., 2024a), and TacoDepth (Wang et al., 2025).
+We consider five recent radar-camera depth estimation baselines: RadarNet [singh2023depth], SparseBeatsDense [li2023sparse], GET-UP [sun2024get], RadarCam-Depth [li2024radarcam], and TacoDepth [wang2025tacodepth].
 Notably, both RadarCam-Depth and TacoDepth also take scaleless MDE predictions as input, together with radar points, to predict metric depth.
 Datasets.
-We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Sec.
-C for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
+We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Supp.
+Mat.
+for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
 4.1 Main Results
@@
 We qualitatively compare POLAR against baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (Fig.
-4).
+3).
 POLAR more accurately predicts metric scale and exhibits fewer global misalignments.
@@
 With a learned polynomial transformation, POLAR corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
-While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure.
-As shown in the top-left highlighted region of Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths.
+RadarCam-Depth predicts a per-pixel scale map; the large number of degrees of freedom can make it challenging to preserve initially correct MDE depth structure.
+As shown in the top-left of Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths.
 Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts.
-This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the construction crane arms in Image B.
-GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
-As the polynomial degree—selected as a hyperparameter—increases, the maximum possible number of inflection points also grows, giving POLAR greater expressive power to correct misalignments between the many scene elements in Fig.
-3, including cars, trucks, trees, buildings, and traffic lights.
+This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the crane arms in Image B.
+GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
+As the degree of the polynomial increases, the maximum possible number of inflection points also grows, giving POLAR greater expressive power to correct misalignments between the many elements of the scene in Fig.
+4, including cars, trucks, trees, buildings, and traffic lights.
 This is reflected in the successively improved error maps.
 See Sec.
-A for quantitative evaluations.
+5 for quantitative evaluations.
+Method Train Time / Epoch
+Lin [lin2020depth] 89.25
+RadarNet [singh2023depth] 101.50
+SparseBeatsDense [li2023sparse] 63.96
+GET-UP [sun2024get] 249.57
+RadarCam-Depth [li2024radarcam] 86.38
+POLAR (Ours) 33.16
+Table 2: Training times (minutes per epoch) on nuScenes using a single NVIDIA A6000 GPU.
+Method Inference Time GFLOPs
+Lin [lin2020depth] 129.96 550.43
+SparseBeatsDense [li2023sparse] 97.47 532.74
+GET-UP [sun2024get] 445.45 630.99
+RadarCam-Depth [li2024radarcam] 315.64 619.02
+TacoDepth [wang2025tacodepth] 29.30 139.87
+POLAR (Ours) 24.81 89.70
+Table 3: Inference time in milliseconds (ms) and computational cost (GFLOPs) on nuScenes.
 4.2 Computational Efficiency
-Table 2: Training times (minutes per epoch) on nuScenes using an NVIDIA A6000 GPU.
-Method Train Time / Epoch
-Lin (Lin et al., 2020) 89.25
-RadarNet (Singh et al., 2023) 101.50
-SparseBeatsDense (Li et al., 2024b) 63.96
-GET-UP (Sun et al., 2025) 249.57
-RadarCam-Depth (Li et al., 2024a) 86.38
-POLAR (Ours) 33.16
-Table 3: Inference time in milliseconds (ms) and computational cost (GFLOPs) on nuScenes.
-Method Inference Time GFLOPs
-Lin (Lin et al., 2020) 129.96 550.43
-SparseBeatsDense (Li et al., 2024b) 97.47 532.74
-GET-UP (Sun et al., 2025) 445.45 630.99
-RadarCam-Depth (Li et al., 2024a) 315.64 619.02
-TacoDepth (Wang et al., 2025) 29.30 139.87
-POLAR (Ours) 24.81 89.70
 POLAR achieves state-of-the-art training time, inference time (ms), and computational overhead (GFLOPs) compared to all baseline methods, while also achieving state-of-the-art accuracy.
@@
 Taken together, these results establish POLAR as not only the most accurate radar-camera depth estimation method, but also the most practical for real-time deployment where latency is critical.
-5 Discussion
-Limitations.
-POLAR requires tuning the polynomial degree as a hyperparameter (Sec.
-A), which may vary across datasets.
-Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
-Summary.
-We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth.
-Our principled and efficient approach demonstrates that a fundamental insight—shifting from affine scale-and-shift to flexible polynomial transformations—outperforms computationally heavier methods.
-Appendix
-Appendix A Comparative Studies
-Table 4: MDE backbone comparative studies on nuScenes.
-nuScenes
-Method MAE RMSE
-DPT (Ranftl et al., 2021) 5188.2 6884.5
-Depth Anything (Yang et al., 2024) 2404.9 4851.1
-Depth Pro (Bochkovskii et al., 2024) 3835.0 6600.3
-UniDepth (Piccinelli et al., 2025) 2129.8 4887.7
-RadarCam-Depth w/ DPT 1689.7 3948.0
-RadarCam-Depth w/ Depth Anything 1953.6 5107.8
-RadarCam-Depth w/ Depth Pro 3417.1 6462.0
-RadarCam-Depth w/ UniDepth 1872.0 4321.2
-POLAR w/ DPT 1525.6 3745.0
-POLAR w/ Depth Anything 1515.1 3719.4
-POLAR w/ Depth Pro 1627.7 4143.9
-POLAR w/ UniDepth 1407.8 3193.5
-MDE Models.
-We evaluate POLAR using four monocular depth estimation (MDE) backbones: DPT (Ranftl et al., 2021), Depth Anything (Yang et al., 2024), Depth Pro (Bochkovskii et al., 2024), and UniDepth (Piccinelli et al., 2025).
-DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see raw MDE performance in Tabs.
-41010), and thus we process them the same way as scaleless depth.
+5 Comparative Studies
+Polynomial Degree.
 Tab.
-4 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and POLAR.
-Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements.
-In contrast, POLAR consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting.
-For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results.
-Among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 51.2% and over RadarCam-Depth w/ UniDepth by 37.8%.
-Table 5: Sensitivity study of polynomial degree, selected as a hyperparameter.
+4 shows that POLAR with polynomial degree 8 outperforms lower degrees.
+At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
+Ablations.
+As shown in Tab.
+5, replacing learnable prototypes with self-attention on the radar point features, or replacing our feature aggregation method (Eq.
+2) with cross-attention, both degrade performance, demonstrating that prototypes capture relevant patterns within radar point configurations.
+Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
+Finally, learning to directly regress depth from our final fused representation using a decoder head performs substantially worse.
+We hypothesize that this is because the number of degrees of freedom ($H\times W$ pixels) is far larger than the available constraints (sparse radar points) to solve for depth at every pixel.
 nuScenes ZJU
@@
 10 1463.7 3494.5 643.3 1184.5
-Table 6: Ablation studies of architecture and loss components.
+Table 4: Sensitivity study of polynomial degree hyperparameter.
 nuScenes VoD
 Ablated Component MAE RMSE MAE RMSE
+cross-modality att. 2238.8 4817.4 2147.9 4926.1
 learnable prototypes 1615.5 3629.0 1619.3 4162.7
-cross-modality att. 2238.8 4817.4 2147.9 4926.1
+feature aggregation 1454.1 3488.9 1543.1 4070.6
+direct decoding 1968.0 4155.7 1855.9 4608.0
 monotonicity loss 1921.1 4399.5 1924.5 4660.3
-pos. embeddings 1454.1 3488.9 1500.1 3951.8
-no ablations 1407.8 3193.5 629.6 1171.3
-Polynomial Degree.
+no ablations 1407.8 3193.5 1500.1 3951.8
+Table 5: Ablation studies of architecture and loss components.
+6 Discussion
+Limitations.
+POLAR requires tuning the polynomial degree as a hyperparameter (Sec.
+5), which may vary across datasets.
+Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
+Summary.
+We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth.
+Our principled and efficient approach demonstrates that a fundamental insight—shifting from affine scale-and-shift to flexible polynomial transformations—outperforms computationally heavier methods.
+[…]
+Supplementary Material
+nuScenes
+Method MAE RMSE
+DPT [ranftl2021vision] 5188.2 6884.5
+UniDepth [piccinelli2025unidepthv2] 2129.8 4887.7
+Depth Anything [depth_anything_v2] 2404.9 4851.1
+Depth Pro [depth-pro] 3835.0 6600.3
+RadarCam-Depth w/ DPT 1689.7 3948.0
+RadarCam-Depth w/ UniDepth 1872.0 4321.2
+RadarCam-Depth w/ Depth Anything 1953.6 5107.8
+RadarCam-Depth w/ Depth Pro 3417.1 6462.0
+TacoDepth w/ DPT 1492.4 3324.8
+POLAR w/ DPT 1447.2 3304.9
+POLAR w/ UniDepth 1407.8 3193.5
+POLAR w/ Depth Anything 1515.1 3580.0
+POLAR w/ Depth Pro 1461.5 4143.9
+Table 6: MDE backbone comparative studies on nuScenes.
+Appendix A Comparative Studies
+MDE Models.
+We evaluate POLAR using four monocular depth estimation (MDE) backbones: DPT [ranftl2021vision], Depth Anything [depth_anything_v2], Depth Pro [depth-pro], and UniDepth [piccinelli2025unidepthv2].
+DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see raw MDE performance in Tabs.
+6711), and thus we process them the same way as scaleless depth.
 Tab.
-6 quantifies the performance gains achieved with higher-degree polynomials, with degree 8 outperforming lower degrees.
-At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
-Figure 5: To quantify absolute improvement in our qualitative results, we provide the colorbars that were used in all examples in Figs. 346, and 7.
-Ablations.
-We evaluate the impact of key components of POLAR in Tab.
-6.
-Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations.
-Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features.
-Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
+6 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and POLAR.
+Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements.
+In contrast, POLAR consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting.
+For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results.
+Among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 51.2% and over RadarCam-Depth w/ UniDepth by 37.8%.
+For ZJU (see Tab. 7), among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 61.1% and over RadarCam-Depth w/ UniDepth by 54.2%.
+TacoDepth.
+Although TacoDepth experiments with only DPT as its MDE backbone, our experiments show that a different backbone provides further performance gains.
+Nevertheless, POLAR outperforms TacoDepth by over 3% in MAE when both methods use the same DPT backbone.
+A similar trend holds on ZJU: even when both methods use the same DPT backbone, POLAR outperforms TacoDepth by 31.5% in MAE, further confirming that POLAR’s state-of-the-art performance is not due to backbone choice.
+ZJU
+Method MAE RMSE
+DPT [ranftl2021vision] 1885.3 3326.1
+UniDepth [piccinelli2025unidepthv2] 1533.0 3188.4
+Depth Anything [depth_anything_v2] 1943.2 3469.3
+Depth Pro [depth-pro] 1680.2 3144.9
+RadarCam-Depth w/ DPT 1183.5 3229.0
+RadarCam-Depth w/ UniDepth 1152.5 3168.6
+RadarCam-Depth w/ Depth Anything 1724.4 3661.3
+RadarCam-Depth w/ Depth Pro 1490.6 3429.5
+TacoDepth w/ DPT 1032.5 2850.3
+POLAR w/ DPT 707.1 1216.9
+POLAR w/ UniDepth 629.6 1171.3
+POLAR w/ Depth Anything 657.2 1225.4
+POLAR w/ Depth Pro 640.3 1174.8
+Table 7: MDE backbone comparative studies on ZJU.
+Figure 5: To quantify absolute improvement in our qualitative results, we provide the colorbars that were used in all examples in Figs. 6 and 7, as well as Figs. 3 and 4 in the main paper.
 Figure 6: Additional qualitative comparison on VoD.
@@
 For nuScenes, where each image has shape $(900,1600)$, each additional polynomial term incurs an additional computational cost of $0.0043$ GFLOPs, which is a $0.0048\%$ increase.
+Cost breakdown.
+Following the reporting protocol of TacoDepth [wang2025tacodepth], we measure inference time as the latency added on top of the MDE backbone.
+Tab.
+8 shows that, even when considering the combined cost, our method remains state-of-the-art in both inference time and GFLOPs.
 Appendix C Dataset Details
@@
 […]
+nuScenes
+Method Inference Time GFLOPs
+Depth Anything [depth_anything_v2] 28.70 201.03
++ Additional over Depth Anything 315.64 619.02
+RadarCam-Depth w/ Depth Anything 344.34 820.05
++ Additional over Depth Anything 29.30 139.87
+TacoDepth w/ Depth Anything 58.00 340.90
++ Additional over Depth Anything 24.81 89.70
+POLAR w/ Depth Anything 53.51 290.73
+Table 8: Computational cost breakdown for methods that leverage MDE. Inference time is measured in milliseconds (ms).
 Appendix D Full nuScenes Benchmark
-Table 7: Full quantitative results ($m$) on the nuScenes benchmark.
 nuScenes
 Distance Method MAE RMSE
-50m NLSPN (Park et al., 2020) 2790.0 5813.4
-BPNet (Tang et al., 2024) 2407.0 4438.0
-RC-PDA (Long et al., 2021) 2225.0 4156.5
-RC-PDA-HG (Long et al., 2021) 2210.0 4234.0
-BTS (Lee et al., 2019) 1937.0 3885.0
-DORN (Lo & Vandewalle, 2021) 1926.6 4124.8
-RadarNet (Singh et al., 2023) 1727.7 3746.8
-CaFNet (Sun et al., 2024) 1674.0 3674.0
-Lin (Lin et al., 2020) 1598.2 3790.1
-SparseBeatsDense (Li et al., 2024b) 1524.5 3567.3
-RadarCam-Depth (Li et al., 2024a) 1286.1 2964.3
-GET-UP (Sun et al., 2025) 1241.0 2857.0
-TacoDepth (Wang et al., 2025) 1046.8 2487.5
+50m RC-PDA [long2021full] 2225.0 4156.5
+RC-PDA-HG [long2021full] 2210.0 4234.0
+NLSPN [Park2020NonLocalSP] 2185.7 4411.6
+BTS [lee2019big] 1937.0 3885.0
+DORN [lo2021depth] 1926.6 4124.8
+RadarNet [singh2023depth] 1727.7 3746.8
+CaFNet [sun2024cafnet] 1674.0 3674.0
+BPNet [tang2024bilateral] 1618.9 3596.1
+Lin [lin2020depth] 1598.2 3790.1
+SparseBeatsDense [li2023sparse] 1524.5 3567.3
+RadarCam-Depth [li2024radarcam] 1286.1 2964.3
+GET-UP [sun2024get] 1241.0 2857.0
+TacoDepth [wang2025tacodepth] 1046.8 2487.5
 POLAR (Ours) 1014.4 2475.7
-70m NLSPN (Park et al., 2020) 3140.0 6580.6
-RC-PDA (Long et al., 2021) 3326.1 6700.6
-RC-PDA-HG (Long et al., 2021) 3485.6 7002.9
-BTS (Lee et al., 2019) 2346.0 4811.0
-DORN (Lo & Vandewalle, 2021) 2170.0 4532.0
-RadarNet (Singh et al., 2023) 2073.2 4590.7
-CaFNet (Sun et al., 2024) 2010.0 4493.0
-Lin (Lin et al., 2020) 1897.8 4558.7
-SparseBeatsDense (Li et al., 2024b) 1822.9 4303.6
-RadarCam-Depth (Li et al., 2024a) 1587.9 3662.5
-GET-UP (Sun et al., 2025) 1541.0 3657.0
-TacoDepth (Wang et al., 2025) 1347.1 3152.8
+70m RC-PDA [long2021full] 3326.1 6700.6
+RC-PDA-HG [long2021full] 3485.6 7002.9
+BTS [lee2019big] 2346.0 4811.0
+NLSPN [Park2020NonLocalSP] 2260.9 4645.0
+DORN [lo2021depth] 2170.0 4532.0
+RadarNet [singh2023depth] 2073.2 4590.7
+CaFNet [sun2024cafnet] 2010.0 4493.0
+Lin [lin2020depth] 1897.8 4558.7
+SparseBeatsDense [li2023sparse] 1822.9 4303.6
+RadarCam-Depth [li2024radarcam] 1587.9 3662.5
+GET-UP [sun2024get] 1541.0 3657.0
+TacoDepth [wang2025tacodepth] 1347.1 3152.8
 POLAR (Ours) 1286.1 2947.3
-80m NLSPN (Park et al., 2020) 3257.7 6872.4
-RC-PDA (Long et al., 2021) 3721.0 7632.0
-RC-PDA-HG (Long et al., 2021) 3664.0 7775.0
-AdaBins (Bhat et al., 2021) 3541.0 5885.0
-P3Depth (Patil et al., 2022) 3130.0 5838.0
-LapDepth (Song et al., 2021) 2544.0 5151.0
-PnP (Wang et al., 2018) 2496.0 5578.0
-BTS (Lee et al., 2019) 2467.0 5125.0
-DORN (Lo & Vandewalle, 2021) 2432.0 5304.0
-RadarNet (Singh et al., 2023) 2179.3 4898.7
-CaFNet (Sun et al., 2024) 2109.0 4765.0
-Lin (Lin et al., 2020) 1988.4 4841.1
-SparseBeatsDense (Li et al., 2024b) 1927.0 4609.6
-RadarCam-Depth (Li et al., 2024a) 1689.7 3948.0
-GET-UP (Sun et al., 2025) 1632.0 3932.0
-TacoDepth (Wang et al., 2025) 1492.4 3324.8
+80m RC-PDA [long2021full] 3721.0 7632.0
+RC-PDA-HG [long2021full] 3664.0 7775.0
+AdaBins [bhat2021adabins] 3541.0 5885.0
+P3Depth [patil2022p3depth] 3130.0 5838.0
+LapDepth [song2021monocular] 2544.0 5151.0
+PnP [wang2018plug] 2496.0 5578.0
+NLSPN [Park2020NonLocalSP] 2519.3 5033.7
+BTS [lee2019big] 2467.0 5125.0
+DORN [lo2021depth] 2432.0 5304.0
+RadarNet [singh2023depth] 2179.3 4898.7
+CaFNet [sun2024cafnet] 2109.0 4765.0
+Lin [lin2020depth] 1988.4 4841.1
+SparseBeatsDense [li2023sparse] 1927.0 4609.6
+RadarCam-Depth [li2024radarcam] 1689.7 3948.0
+GET-UP [sun2024get] 1632.0 3932.0
+TacoDepth [wang2025tacodepth] 1492.4 3324.8
 POLAR (Ours) 1407.8 3193.5
+Table 9: Full quantitative results ($m$) on the nuScenes benchmark.
 We present the full set of quantitative results on the nuScenes dataset in Tab.
-7.
+9.
 This table includes additional baseline methods that were omitted from the main text for brevity, providing a comprehensive comparison of POLAR against all known existing radar-camera depth estimation methods.
@@
 One potential idea for radar-camera depth estimation is to apply lidar-camera depth completion methods designed to densify sparse depth maps using surrounding context.
-One such method, BPNet (Tang et al., 2024) achieves state-of-the-art performance on the KITTI depth completion benchmark by leveraging bilateral propagation.
+One such method, BPNet [tang2024bilateral] achieves state-of-the-art performance on the KITTI depth completion benchmark by leveraging bilateral propagation.
 However, when applied to radar-camera depth estimation, BPNet performs poorly, as shown in Tab.
-7.
-POLAR outperforms BPNet by 57.9% in MAE and 44.2% in RMSE on nuScenes, highlighting the limitations of directly applying lidar depth completion methods to radar data.
+9.
+POLAR outperforms BPNet by 37.3% in MAE and 31.2% in RMSE on nuScenes, highlighting the limitations of directly applying lidar depth completion methods to radar data.
 The key reason for this underperformance lies in the fundamental differences between lidar and radar point clouds.
 Unlike lidar, radar measurements are orders of magnitude sparser and significantly noisier due to factors such as limited antenna elements, elevation ambiguity (see Fig.
-7), and lower range resolution (Singh et al., 2023).
-Lidar depth completion methods assume a relatively dense and structured input (Xia et al., 2023), leveraging local spatial continuity to propagate depth estimates effectively.
+7), and lower range resolution [singh2023depth].
+Lidar depth completion methods assume a relatively dense and structured input [xia2023quadric], leveraging local spatial continuity to propagate depth estimates effectively.
 In contrast, radar points are too sparse for such methods to infer meaningful local depth relationships, leading to poor depth reconstruction when attempting direct densification.
-Table 8: Error metrics for depth estimation. These metrics compute the error between predicted depth $\hat{d}(x)$ and ground truth depth $d(x)$.
 Metric Definition
@@
 RMSE $\downarrow$ $\Big(\tfrac{1}{|\Omega|}\sum_{x\in\Omega}|\hat{d}(x)-d(x)|^{2}\Big)^{1/2}$
-In addition, we compare against Non-Local Spatial Propagation Network (NLSPN) (Park et al., 2020), which achieves near state-of-the-art performance on the KITTI depth completion benchmark by refining sparse lidar depth with an iterative non-local spatial propagation procedure.
+Table 10: Error metrics for depth estimation. These metrics compute the error between predicted depth $\hat{d}(x)$ and ground truth depth $d(x)$.
+In addition, we compare against Non-Local Spatial Propagation Network (NLSPN) [Park2020NonLocalSP], which achieves near state-of-the-art performance on the KITTI depth completion benchmark by refining sparse lidar depth with an iterative non-local spatial propagation procedure.
 NLSPN predicts an initial depth map along with pixel-wise confidences, then refines it by estimating non-local neighbors and their corresponding affinities to selectively propagate depth information.
@@
 However, again, when applied to radar-camera depth estimation, NLSPN performs poorly, as shown in Tab.
-7.
-POLAR outperforms NLSPN by 59.8% in MAE and 55.4% in RMSE on nuScenes.
+9.
+POLAR outperforms NLSPN by 44.1% in MAE and 36.6% in RMSE on nuScenes.
 Instead of relying on depth completion-style densification, POLAR directly learns a transformation function from radar to metric depth by leveraging polynomial fitting.
@@
 In contrast, lidar points yield a denser, image-aligned projection, which is why accurate 3D scene reconstruction with depth completion methods is possible.
-Appendix F Evaluation Metrics and Implementation Details
+Appendix F Evaluation Metrics and Implementation
+Details
 The evaluation metrics used in our study, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), are formulated in Tab.
-8.
+10.
 Lower values equal better performance for both MAE and RMSE.
 Unless specified otherwise, all reported values are in millimeters (mm).
-We train for 60 epochs using a cosine decay learning rate scheduler with learning rate of $5\times 10^{-5}$, and use weighting terms $\lambda_{L_{1}}=1.0,\lambda_{L_{2}}=0.4,\lambda_{mono}=0.25$ for our loss function.
-Appendix G Additional Comparisons
-Full MDE Comparisons.
-For ZJU (see Tab.
-10), among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 61.1% and over RadarCam-Depth w/ UniDepth by 54.2%.
-For VoD (see Tab.
-10), POLAR w/ UniDepth achieves the best performance in MAE, improving over raw UniDepth predictions by 42.4% and over RadarCam-Depth w/ UniDepth by 32.7%, while POLAR w/ Depth Anything achieves the best performance in RMSE, improving over RadarCam-Depth w/ Depth Anything by 37.6% and over raw Depth Anything predictions (inverted and median scaled) by 10.4%.
-Table 9: MDE backbone comparative studies on ZJU-4DRadarCam (ZJU).
-ZJU
-Method MAE RMSE
-DPT (Ranftl et al., 2021) 1885.3 3326.1
-Depth Anything (Yang et al., 2024) 1943.2 3469.3
-Depth Pro (Bochkovskii et al., 2024) 1680.2 3144.9
-UniDepth (Piccinelli et al., 2025) 1533.0 3188.4
-RadarCam-Depth w/ DPT 1183.5 3229.0
-RadarCam-Depth w/ Depth Anything 1724.4 3661.3
-RadarCam-Depth w/ Depth Pro 1490.6 3429.5
-RadarCam-Depth w/ UniDepth 1152.5 3168.6
-POLAR w/ DPT 707.1 1216.9
-POLAR w/ Depth Anything 657.2 1225.4
-POLAR w/ Depth Pro 640.3 1174.8
-POLAR w/ UniDepth 629.6 1171.3
-Table 10: MDE backbone comparative studies on View-of-Delft.
+We train for 60 epochs using a cosine decay learning rate scheduler with learning rate of $5\times 10^{-5}$, and use weighting terms $\lambda_{1}=1.0,\lambda_{2}=0.4,\lambda_{m}=0.25$ for our loss function.
 View-of-Delft
 Method MAE RMSE
-DPT (Ranftl et al., 2021) 4117.9 5498.9
-Depth Anything (Yang et al., 2024) 3270.5 4411.9
-Depth Pro (Bochkovskii et al., 2024) 3275.9 5936.7
-UniDepth (Piccinelli et al., 2025) 2605.6 5691.0
+DPT [ranftl2021vision] 4117.9 5498.9
+UniDepth [piccinelli2025unidepthv2] 2605.6 5691.0
+Depth Anything [depth_anything_v2] 3270.5 4411.9
+Depth Pro [depth-pro] 3275.9 5936.7
 RadarCam-Depth w/ DPT 4013.5 5911.9
+RadarCam-Depth w/ UniDepth 2227.4 5385.8
 RadarCam-Depth w/ Depth Anything 3103.6 6328.7
 RadarCam-Depth w/ Depth Pro 2843.4 6082.0
-RadarCam-Depth w/ UniDepth 2227.4 5385.8
 POLAR w/ DPT 1891.4 4252.6
+POLAR w/ UniDepth 1500.1 3960.5
 POLAR w/ Depth Anything 1770.5 3951.8
 POLAR w/ Depth Pro 1520.2 3987.2
-POLAR w/ UniDepth 1500.1 3960.5
-Leveraging Radar.
-Tab.
-13 shows that replacing radar points with learnable, dataset-specific points worsens MAE by 28.0% and RMSE by 29.9%, demonstrating that we indeed leverage the radar inputs effectively.
-As additional evidence, Tab.
-12 shows that our method, evaluated zero-shot cross-dataset, achieves comparable or better performance than baselines trained on the target datasets.
-Tab.
-12 shows we are more robust to reduced radar point density at inference, i.e., less performance degradation than the baseline method RadarCam-Depth.
-Table 11: Zero-shot generalization.
+Table 11: MDE backbone comparative studies on View-of-Delft.
 $\rightarrowZJU nuScenes\rightarrowVoD$
@@
 Ours (zero-shot)   1147.9 3109.5 2256.2 4744.2
-Table 12: Reduced radar point density.
+Table 12: Zero-shot generalization.
 RadarCam-Depth Ours
@@
 100% kept / 0% removed   1689.7 3948.0 1407.8 3193.5
-Table 13: Additional comparisons of POLAR vs. learnable dataset-specific points in place of radar points, and regression baselines.
+Table 13: Reduced radar point density.
 MAE RMSE
@@
 Our Performance 1407.8 3193.5
+Table 14: Additional comparisons of POLAR vs. learnable dataset-specific points in place of radar points, and regression baselines.
+Appendix G Additional Comparisons
+Full MDE Comparisons.
+For VoD (see Tab.
+11), POLAR w/ UniDepth achieves the best performance in MAE, improving over raw UniDepth predictions by 42.4% and over RadarCam-Depth w/ UniDepth by 32.7%, while POLAR w/ Depth Anything achieves the best performance in RMSE, improving over RadarCam-Depth w/ Depth Anything by 37.6% and over raw Depth Anything predictions (inverted and median scaled) by 10.4%.
+Leveraging Radar.
+Tab.
+14 shows that replacing radar points with learnable, dataset-specific points worsens MAE by 28.0% and RMSE by 29.9%, demonstrating that we indeed leverage the radar inputs effectively.
+As additional evidence, Tab.
+12 shows that our method, evaluated zero-shot cross-dataset, achieves comparable or better performance than baselines trained on the target datasets.
+Radar Sparsity. Tab. 13 shows we are more robust to reduced radar point density at inference, i.e., less performance degradation than the baseline method RadarCam-Depth.
 Adapters.
 Tab.
-13 shows that recent adapter-based finetuning methods LoRA (Hu et al., 2022) and ViT-Adapter (Chen et al., 2022), even with a post-hoc linear fit to projected radar points, do not outperform us.
+14 shows that recent adapter-based finetuning methods LoRA [hu2022lora] and ViT-Adapter [chen2022vision], even with a post-hoc linear fit to projected radar points, do not outperform our method.
 Regression Baselines.
@@
 We hypothesize that this is due to noise in radar points that can be mitigated by learning (Sec.
-3.2).
+3.2 from the main paper).
 Our learned polynomial fit may, in principle, introduce unwanted inversions of initially correct MDE predictions.
@@
 3.4 and Eqs.
-67), which effectively constrains such inversions.
+6, 7 from the main paper), which effectively constrains such inversions.
 To quantify this, we compute Kendall’s $\tau$ coefficient between predicted and ground-truth depths.
@@
 Proof by Construction.
-Consider the following three (MDE prediction $\hat{d}$, ground truth $d$) pairs from the nuScenes dataset, specifically from the image shown in Fig. 3:
+Consider the following three (MDE prediction $\hat{d}$, ground truth $d$) pairs from the nuScenes dataset, specifically from the image shown in Fig. 4 from the main paper:
 $(\hat{d},d)\in\{(5,7),(10,9),(15,44)\}.$
@@
 From the first two pairs, we obtain:
-$7=5\alpha+\beta,\quad 9=10\alpha+\beta.$
+$7=5\alpha+\beta,\qquad 9=10\alpha+\beta.$
 Subtracting gives $\alpha=0.4$ and $\beta=5$ as the unique solution for these two pairs.